### v1

# Adaptive Federated Learning with Auto-Tuned Clients

Abstract
Federated learning (FL) is a distributed machine learning framework where the global model of a central server is trained via multiple collaborative steps by participating clients without sharing their data. While being a flexible framework, where the distribution of local data, participation rate, and computing power of each client can greatly vary, such flexibility gives rise to many new challenges, especially in the hyperparameter tuning on both the server and the client side. We propose $\Delta$-SGD, a simple step size rule for SGD that enables each client to use its own step size by adapting to the local smoothness of the function each client is optimizing. We provide theoretical and empirical results where the benefit of the client adaptivity is shown in various FL scenarios. In particular, our proposed method achieves TOP-1 accuracy in 73% and TOP-2 accuracy in 100% of the experiments considered without additional tuning.
1 Introduction
Federated learning (FL) is a distributed machine learning framework that allows multiple clients to learn a global model collaboratively. Each client trains the model on its local data and sends only the updated model parameters to a central server for aggregation. Mathematically, FL aims to solve the following optimization problem:
$\min_{x\in\mathbb{R}^{d}}f(x):=\frac{1}{m}\sum_{i=1}^{m}f_{i}(x),$ (1)
where $f_{i}(x):=\mathbb{E}_{z\sim\mathcal{D}_{i}}[F_{i}(x,z)]$ is the loss function of the $i$-th client, and $m$ is the number of clients.
A key property of FL is its flexibility in how different clients participate in the overall training: the number of clients $m$, their participation rates, and computing power available to each client can all vary and change at any point during the overall training procedure. On top of that, each client’s local data is not shared with other clients or the server, resulting in certain inherent data privacy McMahan et al. (2017); Agarwal et al. (2018).
While advantageous, such flexibility and data privacy also introduce a plethora of new challenges, notably: $i)$ how the server aggregates the local information coming from each client, and $ii)$ how to make sure each client meaningfully “learns” using their local data and computing device. The first challenge was partially addressed in Reddi et al. (2021), where adaptive optimization methods, such as Adagrad Duchi et al. (2011) or Adam Kingma and Ba (2014), were utilized in the aggregation step. However, the second challenge remains largely unaddressed as we elaborate below, other than the exception of Wang et al. (2021); Xie et al. (2019) reviewed in Section 2.
Local data of each client is not shared, which intrinsically implies heterogeneity in terms of the size and the distribution $\mathcal{D}_{i}$ of local datasets. That is, $\mathcal{D}_{i}$ differs for each client $i$, as well as the number of samples $z\sim\mathcal{D}_{i}$. Consequently, $f_{i}(x)$ can vastly differ from client to client, making Problem (1) hard to solve (i.e., find an approximate minimizer). Moreover, the sheer amount of local updates is far larger than the number of aggregation steps, given that less communication is desired (i.e., less server-side aggregation) due to the high communication cost—typically 3-4$\times$ orders of magnitude more expensive than local computation—in distributed optimization settings Lan et al. (2020).
Figure 1: (A): CIFAR-10 classification task trained on ResNet-18. (B): MNIST classification task trained on shallow CNN. (C): CIFAR-100 classification task trained on ResNet-50. All experiments use FedAvg McMahan et al. (2017) as the server optimizer.
Extensive fine-tuning of the client optimizers is often required to achieve good performance. For instance, experimental results of the famous FedAvg algorithm were obtained after performing a grid-search of typically 11-13 step sizes of the clients’ SGD (McMahan et al., 2017, Section 3), as SGD (and its variants) are highly sensitive to the step size Moulines and Bach (2011); Toulis and Airoldi (2017); Assran and Rabbat (2020); Kim et al. (2022). Similarly, in Reddi et al. (2021), 6 different client step sizes were grid-searched for different tasks, and not surprisingly, each task requires a different client step size to obtain the best result, regardless of the server-side adaptivity (Reddi et al., 2021, Table 8). Importantly, even these “fine-tunings” are done under the setting that all clients use the same step size, which is sub-optimal, given that $f_{i}$ can be vastly different per client; we analyze this further in Section 2.
Initial examination as motivation. The implication of not properly tuning the client optimizer is highlighted in Figure 1. We plot the progress of test accuracies for different client optimizers, where, for all the other test cases, we intentionally use the same step size rules that were fine-tuned for the task in Figure 1(A). There, we train a ResNet-18 for CIFAR-10 dataset classification within an FL setting, where the best step sizes were used for all algorithms after grid-search; we defer the experimental details to Section 4. Hence, all methods perform reasonably well, although $\Delta$-SGD, our proposed method, achieves noticeably better test accuracy when compared to non-adaptive SGD variants –e.g., a 5% difference in final classification accuracy from SGDM– and comparable final accuracy only with adaptive SGD variants, like Adam and Adagrad.
In Figure 1(B), we train a shallow CNN for MNIST data classification using the same step size rules. MNIST classification is now considered an “easy” task, and therefore SGD with the same constant and decaying step sizes from Figure 1(A) works well. However, with momentum, SGDM exhibits highly oscillating behavior, which results in slow progress and poor final accuracy, especially without decaying the step size. Adaptive optimizers, like Adam and Adagrad, show similar behavior, falling short in achieving good final accuracy, compared to their performance in the case of Figure 1(A).
Similarly, in Figure 1(C), we plot the test accuracy progress for CIFAR-100 classification trained on ResNet-50, again using the same step size rules as before. Contrary to Figure 1(B), SGD with momentum (SGDM) works better than SGD, both with the constant and the decaying step sizes. Adam becomes a “good optimizer” again, but its “sibling”, Adagrad, performs worse than SGDM. On the contrary, our proposed method, $\Delta$-SGD, which we introduce in Section 3, achieves superior performance in all cases without any additional tuning.
The above empirical observations beg answers to important and non-trivial questions in training FL tasks using variants of SGD methods as the client optimizer: Should the momentum be used? Should the step size be decayed? If so, when? Unfortunately, Figure 1 indicates that the answers to these questions highly vary depending on the setting; once the dataset itself or how the dataset is distributed among different clients changes, or once the model architecture changes, the client optimizers have to be properly re-tuned to ensure good performance. Perhaps surprisingly, the same holds for adaptive methods like Adagrad Duchi et al. (2011) and Adam Kingma and Ba (2014).
Our hypothesis and contributions. Our paper takes a stab in this direction: we propose DELTA-SGD (DistributEd LocaliTy Adaptive SGD), a simple adaptive SGD scheme, that does not require any knowledge of the problem constants, such as the $L$-smoothness parameter. We will refer to our algorithm as $\Delta$-SGD in the rest of the text. Our adaptive step size is extended from Malitsky and Mishchenko (2020), where the focus was centralized convex optimization; we defer the readers to Section 3 for details. Our contributions can be summarized as follows:
  • We extend the (centralized) adaptive step size from Malitsky and Mishchenko (2020) to the FL setting. The implication of this adaptive step size in FL is twofold: $i)$ each client can use its own step size, and $ii)$ each client’s step size adapts to the local smoothness of $f_{i}$—hence LocaliTy Adaptive— and can even increase during local iterations. Moreover, thanks to the simplicity of the proposed step size, our method is agnostic to the server optimizer as well as the loss function, and thus can be easily combined with server adaptive methods such as FedAdam Reddi et al. (2021), or methods that use different (regularized) loss functions such as FedProx Li et al. (2020) or MOON Li et al. (2021).
  • We provide convergence analysis of $\Delta$-SGD in a general nonconvex setting (Theorem 1). We also provide the convergence analysis in the convex setting, extending the analysis from Malitsky and Mishchenko (2020) to a distributed one, as well as a simple centralized nonconvex convergence analysis, which was an open question posed in Malitsky and Mishchenko (2020). Due to space constraints, we only state the result of general nonconvex case in Theorem 1, and defer other theorems and proofs to Appendix A.
  • We evaluate our approach on several benchmark datasets and demonstrate that $\Delta$-SGD achieves superior performance compared to other state-of-the-art FL methods. Our experiments show that our method can effectively adapt the client step size to the underlying data distribution and achieve faster convergence of the global model, without any additional tuning. Our approach can help overcome the client step size tuning challenge in FL and enable more efficient and effective collaborative learning in distributed systems.
2 Preliminaries and Related Work
A bit of background on optimization theory. Arguably the most fundamental optimization algorithm for minimizing a function $f(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the gradient descent (GD) Cauchy et al. (1847), which iterates with the step size $\eta_{t}$ as: $x_{t+1}=x_{t}-\eta_{t}\nabla f(x_{t})$. Under the assumption that $f$ is globally $L$-smooth, that is:
$\|\nabla f(x)-\nabla f(y)\|\leqslant L\cdot\|x-y\|\quad\forall x,y,$ (2)
the step size $\eta_{t}=\frac{1}{L}$ is the “optimal” step size for GD, which converges at the rate, which is not improvable Nesterov (2003); Drori and Teboulle (2014), for convex $f$:
$f(x_{t+1})-f(x^{\star})\leqslant\frac{L\|x_{0}-x^{\star}\|^{2}}{2(2t+1)},$ (3)
while $\eta_{t}\in(0,\frac{2}{L})$ defines the “admissible” step size range required to ensure that GD converges for this function class Polyak (1963). Similarly, if $f$ is also (globally) $\mu$-strongly convex, that is:
$f(x)\geqslant f(y)+\langle\nabla f(y),x-y\rangle+\tfrac{\mu}{2}\|x-y\|^{2}~{}~{}\forall x,y,$
then the optimal step size is $\eta_{t}=\frac{2}{\mu+L}$ Nesterov (2003), and the rate improves to a linear one:
$\|x_{t+1}-x^{\star}\|^{2}\leqslant\Big{(}\frac{L-\mu}{L+\mu}\Big{)}^{t}\|x_{0}-x^{\star}\|^{2}.$
Limitations of popular step sizes. In practice, however, the constants $L$ and $\mu$ are rarely known Boyd et al. (2004). As such, there has been a plethora of efforts in the optimization community to come up with a general step size rule or method that does not require a priori knowledge of problem constants Nesterov (2015); Orabona and Pál (2016); Levy et al. (2018), so that an optimization algorithm can work as “plug-and-play.”
This is a well-known research question, with some major lines of work on this direction being: $i)$ line-search Armijo (1966); Paquette and Scheinberg (2020), $ii)$ adaptive learning rate (e.g., Adagrad Duchi et al. (2011) and Adam Kingma and Ba (2014)), and $iii)$ Polyak step size Polyak (1969); Hazan and Kakade (2019), to name a few. Unfortunately, the search for the “perfect” step size is still active Loizou et al. (2021); Defazio et al. (2022); Defazio and Mishchenko (2023); Bernstein et al. (2023); Latafat et al. (2023), as the aforementioned approaches have limitations. For instance, to use line search, solving a sub-routine is often required, inevitably incurring additional evaluations of the function and/or the gradient. To use methods like Adagrad and Adam, the knowledge of the initial distance from the solution set $D$, which is nontrivial to find, is required to ensure good performance Defazio and Mishchenko (2023); see also Section 4. Lastly, Polyak step size requires a sharp knowledge of $f(x^{\star})$ Hazan and Kakade (2019), which is not always known a priori, and is unclear what value to use for an arbitrary loss function and model.
Implications in FL scenarios. On top of that, even with the knowledge of $L$, GD can be slow. Indeed, the global $L$-smoothness condition, as described above, needs to be satisfied for all $x$ and $y$, implying even finite $L$ can be arbitrarily large. This results in a small step size, leading to slow convergence, as seen in (3).
Naturally, the challenge is aggravated in the distributed case. To solve (1), under the assumption that each $f_{i}$ is $L_{i}$-smooth, the step size of the form ${1}/{L_{\text{max}}}$ where ${L_{\text{max}}}:=\max_{i}L_{i}$ is often used Yuan et al. (2016); Scaman et al. (2017); Qu and Li (2019), to ensure convergence Uribe et al. (2020). Yet, one can easily imagine a situation where $L_{i}\gg L_{j}$ for $i\neq j,$ in which case the convergence of $f_{j}(x)$ can be arbitrarily slow by using step size ${1}/{L_{\text{max}}}$.
There are plenty of practical examples where such a scenario could happen. For instance, imagine smartphone photo albums of the members of a fishing club versus a bird-watching club. The albums of these two groups could have vastly different data distributions, one skewed with pictures including fishes and rivers, while the other skewed with birds and forests Wah et al. (2011). On top of that, smartphones from different manufacturers nowadays have distinctly different camera filters applied—recently a big selling point for companies like Apple, introducing technologies like “portrait mode,” and Samsung, introducing 100$\times$ zoom capability, both based on image processing techniques—making even similar pictures to have distinct pixel values, when taken with different smartphones.
Therefore, to ensure that each client learns useful information in the FL setting as in (1), ideally: $i)$ each agent should be able to use its own step size, instead of crude ones like $1/L_{\text{max}}$ for all; $ii)$ the individual step size should be “locally adaptive” to the function $f_{i},$ (i.e., even $1/L_{i}$ can be too crude, analogously to the centralized GD); and $iii)$ the step size should not depend on problem constants like $L_{i}$ an $\mu_{i}$, which are often unknown. In Section 3, we introduce DistributEd LocaliTy Adaptive SGD ($\Delta$-SGD), which satisfies all of the aforementioned desiderata.
Related work on FL. FedAvg McMahan et al. (2017) is the simplest and probably the most famous FL algorithm, which takes a (weighted) average of the client parameters in the aggregation step after clients perform some local iterations. While FedAvg has achieved great results empirically, it has been observed recently that it can suffer from model drift Praneeth Karimireddy et al. (2019) induced by the heterogeneity Hsu et al. (2019). At the core, simple averaging is too crude of an operation. As such, FedAvg uses a weighted average heuristic, where the weights are based on the number of data $n_{i}$ for each client $i$ McMahan et al. (2017).
The authors in Reddi et al. (2021) took this idea one step further: it showed that FedAvg can be seen as a special case where both the clients and the server use the SGD optimizer, with a server learning rate of 1. To handle the heterogeneity and the model drift, they proposed to use adaptive optimization scheme at the server, resulting in, e.g., FedAdam or FedYogi, which respectively use Adam Kingma and Ba (2014) and Yogi Zaheer et al. (2018) as the server optimizer, and can improve upon FedAvg in various scenarios. Note that this fundamentally differs from our approach, where we propose a client adaptive optimizer that can be easily combined with sophisticated server-side aggregation methods like FedAdam.
Perhaps the closest works in spirit to our work are AdaAlter Xie et al. (2019) and locality adaptivity in FL Wang et al. (2021). Both works utilize an adaptive client optimizer, similar to Adagrad Duchi et al. (2011), with either delayed or restarted updates on the adaptive variables (e.g., pre-conditioners and momentum buffers). However, as noted in Reddi et al. (2021), AdaAlter incurs twice the memory compared to FedAvg, as AdaAlter requires communicating both the model parameters as well as the “client accumulators,” which is of the same dimension as model parameters. Wang et al. (2021) overcomes this issue by simply restarting the client optimizer states, but requires complicated local and global correction step to achieve good performance. On the contrary, $\Delta$-SGD changes virtually nothing compared to FedAvg other than the simple step size computation, and yet still achieves impressive performance; see Section 4.
Other approaches have handled the heterogeneity by changing the loss function. FedProx Li et al. (2020) adds an $\ell_{2}$-norm proximal term to the individual loss functions such that the local parameters do not diverge too much from the previously communicated global model. Similarly, MOON Li et al. (2021) uses the cross-entropy loss between the current and previous models. As mentioned in the introduction, our proposed method can be seamlessly combined with these approaches.
Another line of works attempts to handle heterogeneity by utilizing control-variates Praneeth Karimireddy et al. (2019); Liang et al. (2019); Karimireddy et al. (2020), a similar idea to variance reduction from single-objective optimization Johnson and Zhang (2013); Gower et al. (2020). While theoretically appealing, these methods either require full gradient computation periodically or are not usable when not all the clients participate, as the clients must maintain a state throughout all rounds.
3 DELTA($\Delta$)-SGD: Distributed locality adaptive SGD
In Malitsky and Mishchenko (2020), a novel step size rule for (centralized) GD was introduced, which adapts to the minimization function’s local geometry. The proposed step size is clever and strikingly simple:
$\eta_{t}=\min\Big{\{}\tfrac{\|x_{t}-x_{t-1}\|}{2\|\nabla f(x_{t})-\nabla f(x_{t-1})\|},\sqrt{1+\theta_{t-1}}\eta_{t-1}\Big{\}},\quad\theta_{t-1}=\eta_{t-1}/\eta_{t-2}.$ (4)
The step size in (4) adapts to the local smoothness of the iterates, that is:
$\|\nabla f(x_{t})-\nabla f(x_{t-1})\|\leqslant L_{t}\cdot\|x_{t}-x_{t-1}\|,\quad\forall t=1,2,\dots$ (5)
and hence can increase during iterations, while the second condition in (4) ensures that $\eta_{t}$ does not increase arbitrarily. By definition, $L_{t}$ in (5) is never bigger than the global $L$ in (2), resulting in faster convergence in practice (Malitsky and Mishchenko, 2020, Section 4). Note that (4) is vastly different from some standard adaptive methods like Adagrad Duchi et al. (2011), which can only decrease the learning rate, as they accumulate non-negative quantities in the denominator of the step size. As a result, the scalar multiplied to their “adaptive” learning rates play a crucial role and requires proper tuning, similarly to vanilla SGD, as evident in Figure 1; see also Table 1 in Section 4.
We now introduce DELTA($\Delta$)-SGD: DistributEd LocaliTy Adaptive SGD, where we extend (4) to the FL setting, with the inclusion of stochasticity and local iterations, as summarized in Algorithm 1. For brevity, we use ~$\tilde{\nabla}f_{i}(x)=\frac{1}{|\mathcal{B}|}\sum\limits_{z\in\mathcal{B}}F_{i}(x,z)$ to denote the stochastic gradients with batch size $|\mathcal{B}|=b.$
Algorithm 1 DELTA($\Delta$)-SGD: DistributEd LocaliTy Adaptive SGD
1:input: $x_{0}\in\mathbb{R}^{d}$, $\eta_{0},\theta_{0},\gamma>0$, and $p\in(0,1)$.
2:for each round $t=0,1,\ldots,T{-}1$ do
3:     sample a subset $\mathcal{S}_{t}$ of clients with size $|\mathcal{S}_{t}|=p\cdot m$
4:     for each machine in parallel for $i\in\mathcal{S}_{t}$ do
5:         Set $x_{t,0}^{i}=x_{t}$
6:         Set $\eta_{t,0}^{i}=\eta_{0}$   and   $\theta_{t,0}^{i}=\theta_{0}$
7:         for local step $k\in[K]$ do
8:              $x_{t,k}^{i}=x_{t,k-1}^{i}-\eta_{t,k-1}^{i}\tilde{\nabla}f_{i}(x_{t,k-1}^{i})$triangleright update local parameter with $\Delta$-SGD
9:               $\eta_{t,k}^{i}=\min\Big{\{}\frac{\gamma\|x_{t,k}^{i}-x_{t,{k-1}}^{i}\|}{2\|\tilde{\nabla}f_{i}(x_{t,k}^{i})-\tilde{\nabla}f_{i}(x_{t,{k-1}}^{i})\|},\sqrt{1+\theta_{t,k-1}^{i}}\eta_{t,k-1}^{i}\Big{\}}$
10:              $\theta_{t,k}^{i}=\eta_{t,k}^{i}/\eta_{t,k-1}^{i}$triangleright (line 10 & 11) update locality adaptive step size
11:         end for
12:     end for
13:     $x_{t+1}=\frac{1}{|\mathcal{S}_{t}|}\sum\limits_{i\in\mathcal{S}_{t}}x_{t,K}^{i}$triangleright server-side aggregation
14:end for
15:return $x_{T}$
We make a few remarks about Algorithm 1. First, the input $\eta_{0}>0$ can be quite arbitrary, as it can be corrected, per client level, in the 1st local iteration (line 9); similarly for $\theta_{0}>0$ (line 10). Second, as the step size $\eta_{t,k}^{i}$ requires two successive stochastic gradients, there could be two options: using $i)$ the same batch or $ii)$ a new batch; the second option will incur additional memory requirement, so we use the first. [^1] This distinction was already analyzed in the centralized setting, and it was reported that using the same batch performed better in practice Malitsky and Mishchenko (2020). Third, we include the “amplifier” $\gamma$ to the first condition of step size (line 9), but this is only needed for Theorem 1.222In practice, we did not change from the default value $\gamma=2$ in the original implementation in https://github.com/ymalitsky/adaptive_GD/blob/master/pytorch/optimizer.py. fourth, after local steps are done, $\eta_{0}$ is re-used (line 6); we could potentially use a more sophisticated approach and use a similar adaptive step size to line 9, but we leave this for future work. Last, we present simple averaging as the server-side aggregation for simplicity (line 13), but more sophisticated methods like FedAdam Reddi et al. (2021) can be used instead.
3.1 Convergence Analysis
Technical challenges. Before we present the convergence analysis, let us elaborate on the non-trivial technical challenges of analyzing Algorithm 1. The main difference of analyzing Algorithm 1 compared to other distributed/decentralized optimization theory is that the step size $\eta_{t,k}^{i}$ not only depends on the round $t$ and local steps $k$ (which is a challenge shared by other analyses), but also on $i$, the client. To deal with the client-dependent step size, we require a slightly non-standard assumption on the dissimilarity between $f$ and $f_{i}$, as shown below.
Assumption 1.
There exist nonnegative constants $\sigma,\rho,$ and $G$ such that for all $i\in[M]$ and $x\in\mathbb{R}^{d}$,
$\mathbb{E}\|\nabla F_{i}(x,z)-\nabla f_{i}(x)\|$ $\leqslant\sigma^{2},$ (bounded variance) (1a)
$\|\nabla f_{i}(x)\|$ $\leqslant G,$ (bounded gradient) (1b)
$\|\nabla f_{i}(x)-\nabla f(x)\|^{2}$ $\leqslant\rho\|\nabla f(x)\|^{2}.$ (strong growth of dissimilarity) (1c)
Assumption 1a has been standard in stochastic optimization literature Moulines and Bach (2011); Ghadimi and Lan (2013); Stich (2018); Khaled et al. (2020); recently, this assumption has been relaxed to weaker ones such as weak growth assumption Bottou et al. (2018) or expected smoothness Gower et al. (2021), but this is out of scope of this work. Assumption 1b is fairly standard in nonconvex optimization Zaheer et al. (2018); Ward et al. (2020), and often used in FL setting Xie et al. (2019, 2019); Reddi et al. (2021). Assumption 1c is reminiscent of the strong growth assumption in stochastic optimization literature Schmidt and Roux (2013), which is still used in recent works Cevher and Vũ (2019); Vaswani et al. (2019); Vaswani et al. (2019). To the best of our knowledge, this is the first theoretical analysis of the FL setting where clients can use their step size. We now present the main convergence theorem of Algorithm 1.
Theorem 1.
Let Assumption 1 hold, with $\rho=\mathcal{O}(1)$. Further, suppose that $\gamma=\mathcal{O}(\frac{1}{K\sqrt{T}})$, and $\eta_{0}=\mathcal{O}(\gamma)$. Then, the following property holds for Algorithm 1, for $T$ sufficiently large:
$\frac{1}{T}\sum_{t=0}^{T-1}\,\mathbb{E}\left\|\nabla f\left(x_{t}\right)\right\|^{2}$ $\leqslant\mathcal{O}\left(\frac{\Psi_{1}}{\sqrt{T}}\right)+\mathcal{O}\left(\frac{\tilde{L}^{2}\Psi_{2}}{T}\right)+\mathcal{O}\left(\frac{\tilde{L}^{3}\Psi_{2}}{\sqrt{T^{3}}}\right),$
where $\Psi_{1}=\max\left\{\frac{\sigma^{2}}{b},f(x_{0})-f(x^{\star})\right\}$ and $\Psi_{2}=\left(\frac{\sigma^{2}}{b}+G^{2}\right)$ are global constants, with $b=|\mathcal{B}|$ being the batch size; $\tilde{L}$ is a constant at most the average of local smoothness, i.e., $\max_{t}\frac{1}{m}\sum_{i=1}^{m}L_{i,t},$ where $L_{i,t}$ the local smoothness of $f_{i}$ at round $t$.
The convergence result in Theorem 1 implies a sublinear convergence to an $\varepsilon$-first order stationary point with at least $T=\mathcal{O}(\varepsilon^{-2})$ communication rounds. We emphasize again that the conditions on $\gamma$ and $\eta_{0}$ are only required for our theory. [^3] In all experiments in Section 4, we use the default setting $\gamma=2$, and $\eta_{0}=0.2$, without additional tuning. Importantly, we did not assume $f_{i}$ is $L$-smooth (or even $L_{i}$-smooth), a standard assumption in nonconvex optimization and FL literature Ward et al. (2020); Koloskova et al. (2020); Li et al. (2020); Reddi et al. (2021); instead, we can obtain a smaller quantity, $\tilde{L}$, through our analysis in a similar vein to (5); for space limitation, we defer the details and the proof to Appendix A.
4 Experimental Setup and Results
Datasets and models. To test the performance of different client optimizers in various settings, we evaluate on four datasets commonly used in FL scenarios: MNIST, FMNIST, CIFAR-10, and CIFAR-100 Krizhevsky et al. (2009). For MNIST and FMNIST, we train a shallow CNN with two convolutional and two fully-connected layers, followed by dropout and ReLU activations. For CIFAR-10, we train a ResNet-18 He et al. (2016). For CIFAR-100, we train both a ResNet-18 and a ResNet-50 to study the effect of changing the model architecture.
For each dataset, we create a federated version by randomly partitioning the training data among $100$ clients, with each client getting $500$ examples. [^4] All the datasets we consider have $50,000$ training samples, leading to $500$ samples distributed per client. To control the level of “non-iidness”, we apply latent Dirichlet allocation (LDA) over the labels following Hsu et al. (2019), where each client has an associated multinomial distribution over the labels from which its examples are drawn. That is, the local training examples of each client has been drawn from a categorical distribution over $N$ classes, parameterized by $\mathbf{q}$, such that $\mathbf{q}\sim\text{Dir}(\alpha\mathbf{p}),$ where $\mathbf{p}$ is a prior distribution over $N$ classes, with $\alpha>0$ being the concentration parameter. We vary $\alpha\in\{0.01,0.1,1\}$, where bigger $\alpha$ indicates settings closer to i.i.d. scenarios. The class distribution of different $\alpha$’s is visualized in Figure 5 in Appendix B.
FL setup and optimizers. For all cases, we fix the number of clients to be $100,$ and randomly sample $10\%$ among the $100$ clients. Similarly to Reddi et al. (2021); Li et al. (2020), we perform $E$ local epochs of training over each client’s dataset, and we utilize mini-batch gradients of size $b=64$, leading to $K\approx\lfloor\tfrac{E\cdot 500}{64}\rfloor$ local gradient steps; we use $E=1$ for all settings. For client optimizers, we compare stochastic gradient descent (SGD), SGD with momentum (SGDM), adaptive methods including Adam Kingma and Ba (2014), Adagrad Duchi et al. (2011), SGD with stochastic Polyak step size (SPS) Loizou et al. (2021), and our proposed method: $\Delta$-SGD in Algorithm 1. As our focus is on client adaptivity, we only present the results using simple FedAvg McMahan et al. (2017) as the server optimizer.
[…]
Implementation. We use Pytorch Paszke et al. (2019) to implement all experiments and run on eight P100 GPUs. Our implementation will be available online after the anonymous review process for reproducibility.
Figure 2: The effect of stronger heterogeneity on different client optimizers, induced by the Dirichlet concentration parameter $\alpha$.
4.1 Results
Results of the experiments detailed in the previous subsection are presented in Table 1. Below, we make comprehensive remarks, depending on: $i)$ the level of heterogeneity induced by different $\alpha$’s of the Dirichlet concentration parameter, $ii)$ the effect of changing the model architecture, and $iii)$ the effect of changing the dataset, and how these affect the performance of different client optimizers.
Non-iidness Optimizer Dataset / Model
$\text{Dir}(\alpha\cdot\mathbf{p})$ MNIST FMNIST CIFAR-10 CIFAR-100 CIFAR-100
CNN CNN ResNet-18 ResNet-18 ResNet-50
$\alpha=1$ SGD 98.3${}_{\downarrow(0.2)}$ 86.5${}_{\downarrow(0.8)}$ 87.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 57.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.2)}}$ 53.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.8)}}$
SGD ($\downarrow$) 97.8${}_{\downarrow(0.7)}$ 86.3${}_{\downarrow(1.0)}$ 87.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.0)}}$ 61.9${}_{\downarrow(0.0)}$ 60.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.9)}}$
SGDM 98.5${}_{\downarrow(0.0)}$ 85.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 88.7${}_{\downarrow(1.1)}$ 58.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.1)}}$ 60.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.3)}}$
SGDM ($\downarrow$) 98.4${}_{\downarrow(0.1)}$ 87.2${}_{\downarrow(0.1)}$ 89.3${}_{\downarrow(0.5)}$ 61.4${}_{\downarrow(0.5)}$ 63.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.5)}}$
Adam 94.7${}_{{\color[rgb]{0.96,0.29,0.54}{\color[rgb]{0.96,0.29,0.54}\downarrow(3.8)}}}$ 71.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.5)}}$ 89.4${}_{\downarrow(0.4)}$ 55.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.3)}}$ 61.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.4)}}$
Adagrad 64.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(34.2)}}$ 45.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(41.8)}}$ 86.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.2)}}$ 53.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.4)}}$ 51.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.9)}}$
SPS 10.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(88.4)}}$ 85.9${}_{\downarrow(1.4)}$ 82.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$ 1.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(60.9)}}$ 50.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.8)}}$
$\Delta$-SGD 98.4${}_{\downarrow(0.1)}$ 87.3${}_{\downarrow(0.0)}$ 89.8${}_{\downarrow(0.0)}$ 61.5${}_{\downarrow(0.4)}$ 65.8${}_{\downarrow(0.0)}$
$\alpha=0.1$ SGD 98.1${}_{\downarrow(0.0)}$ 83.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.8)}}$ 72.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.9)}}$ 54.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.7)}}$ 44.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.9)}}$
SGD ($\downarrow$) 98.0${}_{\downarrow(0.1)}$ 84.7${}_{\downarrow(1.7)}$ 78.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.6)}}$ 59.3${}_{\downarrow(1.8)}$ 48.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.4)}}$
SGDM 97.6${}_{\downarrow(0.5)}$ 83.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.8)}}$ 79.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.4)}}$ 58.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.3)}}$ 52.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.8)}}$
SGDM ($\downarrow$) 98.0${}_{\downarrow(0.1)}$ 86.1${}_{\downarrow(0.3)}$ 77.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$ 60.4${}_{\downarrow(0.7)}$ 52.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.3)}}$
Adam 96.4${}_{\downarrow(1.7)}$ 80.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.0)}}$ 85.0${}_{\downarrow(0.0)}$ 55.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.7)}}$ 58.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.9)}}$
Adagrad 89.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.2)}}$ 46.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(40.1)}}$ 84.1${}_{\downarrow(0.9)}$ 49.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.5)}}$ 48.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.1)}}$
SPS 96.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 85.0${}_{\downarrow(1.4)}$ 70.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.7)}}$ 42.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(18.9)}}$ 42.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(21.9)}}$
$\Delta$-SGD 98.1${}_{\downarrow(0.0)}$ 86.4${}_{\downarrow(0.0)}$ 84.5${}_{\downarrow(0.5)}$ 61.1${}_{\downarrow(0.0)}$ 64.1${}_{\downarrow(0.0)}$
$\alpha=0.01$ SGD 96.8${}_{\downarrow(0.7)}$ 79.0${}_{\downarrow(1.2)}$ 22.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.3)}}$ 30.5${}_{\downarrow(1.3)}$ 24.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$
SGD ($\downarrow$) 97.2${}_{\downarrow(0.3)}$ 79.3${}_{\downarrow(0.9)}$ 33.9${}_{\downarrow(0.0)}$ 30.3${}_{\downarrow(1.5)}$ 24.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.8)}}$
SGDM 77.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.6)}}$ 75.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.5)}}$ 28.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.5)}}$ 24.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.0)}}$ 22.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.4)}}$
SGDM ($\downarrow$) 94.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.5)}}$ 79.5${}_{\downarrow(0.7)}$ 29.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.9)}}$ 20.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(10.9)}}$ 14.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.7)}}$
Adam 80.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.7)}}$ 60.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.6)}}$ 22.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.8)}}$ 18.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.6)}}$ 22.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.8)}}$
Adagrad 72.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(25.1)}}$ 45.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(34.3)}}$ 12.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(21.4)}}$ 25.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.0)}}$ 22.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.2)}}$
SPS 69.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(27.8)}}$ 44.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(36.2)}}$ 21.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.4)}}$ 22.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.8)}}$ 17.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.0)}}$
$\Delta$-SGD 97.5${}_{\downarrow(0.0)}$ 80.2${}_{\downarrow(0.0)}$ 31.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.3)}}$ 31.8${}_{\downarrow(0.0)}$ 31.4${}_{\downarrow(0.0)}$
Table 1: Experimental results based on the settings detailed in Section 4. Performance difference within 0.5% of the best result for each task are shown in bold. Subscripts${}_{\downarrow(x.x)}$ is the performance difference from the best result and is highlighted in pink when the difference is bigger than 2%. The symbol ($\downarrow$) appended to SGD and SGDM indicates step-wise learning rate decay, where the step sizes are divided by $10$ after 50%, and another by $10$ after 75% of the total rounds.
Changing the level of non-iidness. We first investigate how the performance of different client optimizers degrade in increasing degrees of heterogeneity by varying the concentration parameter $\alpha\in\{1,0.1,0.01\}$ multiplied to the prior of Dirichlet distribution, following Hsu et al. (2019).
Three illustrative cases are visualized in Figure 2. We remind that the hyperparameters are tuned for the case of CIFAR-10 trained with ResNet-18, with $\alpha=0.1$ (Figure 2(A)). For this task, with $\alpha=1$ (i.e., closer to iid), all methods perform better than $\alpha=0.1,$ as expected. With $\alpha=0.01,$ which is highly non-iid (c.f., Figure 5 in Appendix B), we see a significant drop in performance for all methods. SGD with LR decay and $\Delta$-SGD perform the best, while adaptive methods like Adam ($85\%\rightarrow 22.1\%$) and Adagrad ($84.1\%\rightarrow 12.5\%$) degrade noticeably more than other methods.
For FMNIST trained with CNN (Figure 2(B)), which can be considered as an easier problem compared to CIFAR-10 classification, the performance degradation with varying $\alpha$’s is much milder. Interestingly, adaptive methods like Adam, Adagrad, and SPS perform much worse than other methods, indicating their inflexibility when applied to diverse datasets and model architectures, which often is worse than simple SGD-based methods; we further investigate this phenomenon in more detail in the next remark.
A similar trend can be seen in MNIST classification, where SGD-based methods work relatively well, except for SGDM with $\alpha=0.01$: without the LR decay, SGDM only achieves around 78% accuracy, while SGD without LR decay still achieves over $96\%$ accuracy (c.f., Figure 1(A)). Even with the LR decay, SGDM achieves noticeably worse performance. While this may seem like an “easy fix” in hindsight, it indicates that in a highly non-iid setting, one should be extremely careful in tuning SGD-based methods, even for easy datasets like MNIST. Interestingly, SPS performs particularly poorly, merely achieving 10% for MNIST with $\alpha=1,$ while achieving 96% for “harder” case of $\alpha=0.1,$ and again dropping to 69.7% with $\alpha=0.01$; a possible explanation is, with simple CNN, the model is not over-parameterized enough to have $f_{i}^{\star}=0,$ the default estimate we use on which the performance of SPS crucially rely Hazan and Kakade (2019). Adagrad similarly performs quite poorly.
Lastly, in Figure 2(C), results for CIFAR-100 classification trained with ResNet-50 are illustrated. $\Delta$-SGD exhibits superior performance in all cases of $\alpha$. Unlike MNIST and FMNIST, Adam enjoys the second ($\alpha=0.1$) or the third ($\alpha=1$) best performance in this task, complicating how one should “tune” Adam for the task at hand. Other methods, including SGD with and without momentum/LR decay, Adagrad, and SPS, perform much worse than $\Delta$-SGD.
Changing the model architecture. For this remark, let us focus on CIFAR-100 trained on ResNet-18 versus on ResNet-50, with $\alpha=0.1$, illustrated in Figure 3(A). SGD and SGDM (both with and without LR decay), Adagrad, and SPS perform worse using ResNet-50 than ResNet-18. This is a counter-intuitive behavior, as one would expect to get better accuracy by using a more powerful model. $\Delta$-SGD is an exception: without any additional tuning, $\Delta$-SGD can improve its performance. Adam also improves similarly, but the achieved accuracy is significantly worse than that of $\Delta$-SGD. While it is true that the performance of $\Delta$-SGD also slightly degrades for $\alpha=0.01,$ it still performs better than the other methods.
Changing the dataset. We now focus on cases where the dataset changes, but the model architecture remains the same. We mainly consider two cases, illustrated in Figure 3(B) and (C): When CNN is trained for classifying MNIST versus FMNIST, and when ResNet-18 is trained for CIFAR-10 versus CIFAR-100. Aggravating the complexity of fine-tuning SGD(M) for MNIST, one can observe a similar trend in CIFAR-10 trained with ResNet-18. In that case, $\Delta$-SGD does not achieve the best test accuracy (although it is the second best with a pretty big margin with the rest), while SGD with LR decay does. However, without LR decay, the accuracy achieved by SGD drops more than 11%.
Interestingly, for the same setting, when we change the dataset from CIFAR-10 to CIFAR-100, SGD without LR decay (which is not one of the best methods for the CIFAR-10 case) achieves the second-best performance (after $\Delta$-SGD). On the other hand, SGDM with LR decay, which achieves the 3rd best performance for CIFAR-10 achieves very poor performance (more than 10% less than the best) for CIFAR-100. Finally, SGD with decay, which performs the best in the CIFAR-10 case, turns out to be the 3rd best algorithm when we change the dataset to CIFAR-100. Again, adaptive methods like Adam, Adagrad, and SPS perform quite poorly in this case.
Figure 3: The effect of changing the dataset and the model architecture on different client optimizers. (A): CIFAR-100 trained with ResNet-18 versus Resnet-50 ($\alpha=0.1$), (B): MNIST versus FMNIST trained with CNN ($\alpha=0.01$), (C): CIFAR-10 versus CIFAR-100 trained with ResNet-18 ($\alpha=0.01$).
5 Conclusion
In this work, we propose $\Delta$-SGD, a distributed SGD scheme equipped with an adaptive step size that enables each client to use its step size and adapts to the local smoothness of the function each client is optimizing. We prove the convergence of $\Delta$-SGD in the general nonconvex setting and present extensive empirical results, where the superiority of $\Delta$-SGD is shown in various scenarios without any tuning. For future works, extending $\Delta$-SGD to a coordinate-wise step size in spirit of Duchi et al. (2011); Kingma and Ba (2014), as well as enabling asynchronous updates Assran et al. (2020); Toghani et al. (2022); Nguyen et al. (2022) could be interesting directions.
Broader Impacts and Limitations. We believe our proposed method, $\Delta$-SGD, can potentially impact reducing the amount of tuning required in training FL models. By now, it is widely believed that FL systems do not “dodge the bullet” of high carbon emission in AI/ML: as noted in a recent study Yousefpour et al. (2023), “compute on client devices, and the communication between the clients and the server are responsible for the majority of FL’s overall carbon emissions (97%).” The lack of principled hyperparameter tuning only exacerbates this problem. Based on the results presented in Section 4, $\Delta$-SGD achieves superior performance while robust to different settings.
In terms of limitations, our assumption on Theorem 1, especially Assumption 1c, as well as the conditions on $\gamma$ and $\eta_{0}$ are quite strong, and does not truly capture the strength of $\Delta$-SGD. Moreover, due to non-optimal implementation, $\Delta$-SGD takes about 1.3×1.3$\times$1.3 × more wall-clock time compared to SGD, which can be improved with more efficient implementation.
Supplementary Materials for
“Adaptive Federated Learning with Auto-Tuned Clients”
In this appendix, we provide all missing proofs that were not present in the main text, as well as additional plots and experiments. The appendix is organized as follows.
  • In Appendix A, we provide all the missing proofs. In particular:
    • In Section A.1, we provide the proof of Algorithm 1, under Assumption 1.
    • In Section A.2, we provide the proof of Algorithm, 1, with some additional assumptions, including that $f_{i}$ is convex for all $i\in[m]$.
    • In Section A.3, we provide a simple proof in the centralized nonconvex setting, which was an open question posed in Malitsky and Mishchenko (2020).
  • In Appendix B, we provide additional experiments, as well as miscellaneous plots that were missing in the main text due to the space constraints. Specifically:
    • In Section B.1, we provide some perspectives on the ease of tuning of $\Delta$-SGD in Algorithm 1.
    • In Section B.2, we provide illustration of the different level of heterogeneity induced by the Dirichlet concentration parameter $\alpha$.
    • In Section B.3, we provide some perspectives on how $\Delta$-SGD performs using different number of local iterations.
    • In Section B.4, we provide additional experimental results where each client has different number of samples as local dataset.
Appendix A Missing Proofs
We first introduce the following two inequalities which will be used in the following proofs.
$\langle x_{i},x_{j}\rangle$ $\leqslant\frac{1}{2}\|x_{i}\|^{2}+\frac{1}{2}\|x_{j}\|^{2},\quad\text{and}$ (7)
$\Big{\|}\sum_{i=1}^{m}x_{i}\Big{\|}^{2}$ $\leqslant m\sum_{i=1}^{m}\Big{\|}x_{i}\big{\|}^{2},$ (8)
for any set of $m$ vectors $\{x_{i}\}_{i{=}1}^{m}$ such that $x_{i}\in\mathbb{R}^{d}$.
We also provide the following lemma, which establishes $\tilde{L}$-smoothness based on local smoothness, without assuming the global $L$-smoothness.
Lemma 2.
For locally smooth function $f_{i}$, the following holds for the points that Algorithm 1 visits:
$f_{i}(x_{t})\leqslant f_{i}(x_{t-1})+\langle\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$ (9)
Proof.
By the definition of $\eta_{t}^{i},$ we have
$\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1})\|\leqslant\frac{1}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|$
Thus, we have
$\langle\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle$ $\leqslant\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1})\|\cdot\|x_{t}-x_{t-1}\|$
$\leqslant\frac{1}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
Let $g(k)=f_{i}(x_{t-1}+k(x_{t}-x_{t-1})).$ Then, we have
$g^{\prime}(k)$ $=\langle\nabla f_{i}(x_{t-1}+k(x_{t}-x_{t-1})),x_{t}-x_{t-1}\rangle$
$g^{\prime}(k)-g^{\prime}(0)$ $=\langle\nabla f_{i}(x_{t-1}+k(x_{t}-x_{t-1})-\nabla f_{i}(x_{t}),\frac{k}{k}(x_{t}-x{t-1})\rangle$
$\leqslant\frac{k}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
Using the Mean Value Theorem, we have
[…]
$\leqslant g(0)+g^{\prime}(0)+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}$
$\leqslant f_{i}(x_{t-1})+\langle\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
We also establish the averaged version of local $\tilde{L}$-smoothness. First, we re-write the statement of Lemma 2 as follows:
$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}\leqslant\tilde{L}_{i}\|x-y\|^{2},$ (10)
where $\tilde{L}_{i}$ is $\max_{t}\Big{(}\frac{1}{2\eta_{t}^{i}}\Big{)}$ for all $x,y\in\text{conv}\{x_{0},x_{1},\dots\}$, and $\text{conv}\{\cdots\}$ denotes the convex hull of the arguments Boyd et al. (2004). Now, given (10),
Lemma 3.
The average of $\tilde{L}$-local smooth functions $f_{i}$ are at least $\tilde{L}$-local smooth, where $\tilde{L}=\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i}$.
Proof.
Starting from (10), we have:
$\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\nabla f_{i}(x)-\frac{1}{m}\sum_{i=1}^{m}\nabla f_{i}(y)\Big{\|}$ $=\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\big{(}\nabla f_{i}(x)-\nabla f_{i}(y)\big{)}\Big{\|}$
$\leqslant\frac{1}{m}\sum_{i=1}^{m}\big{\|}\nabla f_{i}(x)-\nabla f_{i}(y)\big{\|}$
$\leqslant\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i}\|x-y\|$
$=\tilde{L}\|x-y\|.$ (11)
A.1 Proof of Algorithm 1 under Assumption 1
Proof.
Based on Algorithm 1, we use the following notations throughout the proof:
$x_{t,0}^{i}$ $=x_{t},$
$x_{t,k}^{i}$ $=x_{t,k-1}^{i}-\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k-1}^{i}),\quad\forall k\in[K],$
$x_{t+1}$ $=\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}x_{t,k}^{i},$
where we denoted with $x_{t+1}$ as the server parameter at round $t+1$, which is the average of the local parameters $x_{t,k}^{i}$ for all client $\forall~{}i\in[m]$ after running $k$ local iterations for $\forall~{}k\in[K]$.
We start from the smoothness of $f(\cdot)$, which we emphasize again that is obtained via Lemma 3, i.e., as a byproduct of our analysis, and we do not assume that $f_{i}$’s are globally $L$-smooth. Also note that Lemma 3 holds for $\text{conv}\{x_{0},x_{1},\dots\}$, i.e., the average of the local parameters visited by Algorithm 1.
Proceeding, we have
$f(x_{t+1})\leqslant f(x_{t})-\Big{\langle}\nabla f(x_{t}),\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\rangle}+\frac{\tilde{L}}{2}\Big{\|}\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\|}^{2}.$
Taking expectation from the expression above, we get
$\mathbb{E}f(x_{t+1})$
$\leqslant f(x_{t})-\mathbb{E}\Big{\langle}\nabla f(x_{t}),\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\rangle}+\frac{\tilde{L}}{2}\mathbb{E}\Big{\|}\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\|}^{2}$ (12)
$\leqslant f(x_{t})-\mathbb{E}\Big{\langle}\nabla f(x_{t}),\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\nabla f_{i}(x_{t,k}^{i})\Big{\rangle}+\frac{\tilde{L}}{2m}\sum_{i=1}^{m}\,\mathbb{E}\Big{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\|}^{2}$ (13)
$=f(x_{t})-D_{t}\,\|\nabla f(x_{t})\|^{2}-D_{t}\,\mathbb{E}\Big{\langle}\nabla f(x_{t}),\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\frac{\eta_{t,k}^{i}}{D_{t}}\Big{(}\nabla f_{i}(x_{t,k}^{i})-\nabla f(x_{t})\Big{)}\Big{\rangle}$
$\hskip 113.81102pt+\frac{\tilde{L}}{2m}\sum_{i=1}^{m}\,\mathbb{E}\Big{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\|}^{2},$ (14)
where in the equality, we added and subtracted $\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\nabla f(x_{t})$ in the inner product term, and also replaced $D_{t}=\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}$. Then, using (7) and (12)-(14), we have
$\mathbb{E}f(x_{t+1})$ $\leqslant f(x_{t})-D_{t}\,\|\nabla f(x_{t})\|^{2}+\frac{D_{t}}{2}\,\|\nabla f(x_{t})\|^{2}$ (15)
$\quad+\frac{D_{t}}{2}\,\mathbb{E}\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\frac{\eta_{t,k}^{i}}{D_{t}}\Big{(}\nabla f_{i}(x_{t,k}^{i})-\nabla f(x_{t})\Big{)}\Big{\|}^{2}+\frac{\tilde{L}}{2m}\sum_{i=1}^{m}\,\mathbb{E}\Big{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\|}^{2}$
$=f(x_{t})-\frac{D_{t}}{2}\,\|\nabla f(x_{t})\|^{2}+\frac{1}{2D_{t}}\,\mathbb{E}\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\Big{(}\nabla f_{i}(x_{t,k}^{i})-\nabla f(x_{t})\Big{)}\Big{\|}^{2}$
$\quad+\frac{\tilde{L}}{2m}\sum_{i=1}^{m}\,\mathbb{E}\Big{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\|}^{2}$ (16)
$=f(x_{t})-\frac{D_{t}}{2}\,\|\nabla f(x_{t})\|^{2}$
$\quad+\frac{1}{2D_{t}}\,\mathbb{E}\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\Big{(}\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t,k}^{i})+\nabla f_{i}(x_{t})-\nabla f(x_{t})\Big{)}\Big{\|}^{2}$
$\quad+\frac{\tilde{L}}{2m}\sum_{i=1}^{m}\,\mathbb{E}\Big{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\Big{(}\tilde{\nabla}f_{i}(x_{t,k}^{i})-\nabla f_{i}(x_{t,k}^{i})+\nabla f_{i}(x_{t,k}^{i})-\nabla f_{i}(x_{t})$
$\qquad\qquad\qquad\qquad\qquad+\nabla f_{i}(x_{t})-\nabla f(x_{t})+\nabla f(x_{t})\Big{)}\Big{\|}^{2}$ (17)
We now expand the terms in (17), and will bound each seprately, as follows.
$\mathbb{E}f(x_{t+1})$ $\leqslant f(x_{t})-\frac{D_{t}}{2}\,\|\nabla f(x_{t})\|^{2}$
$\quad+\underbrace{\frac{1}{D_{t}}\,\mathbb{E}\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\Big{(}\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t,k}^{i})\Big{)}\Big{\|}^{2}}_{\mathrm{A}_{1}}$
$\quad+\underbrace{\frac{1}{D_{t}}\,\mathbb{E}\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\Big{(}\nabla f_{i}(x_{t})-\nabla f(x_{t})\Big{)}\Big{\|}^{2}}_{\mathrm{A}_{2}}$
$\quad+\underbrace{\frac{2\tilde{L}}{m}\sum_{i=1}^{m}\,\mathbb{E}\Bigg{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\Big{(}\tilde{\nabla}f_{i}(x_{t,k}^{i})-\nabla f_{i}(x_{t,k}^{i})\Big{)}\Bigg{\|}^{2}}_{\mathrm{A}_{3}}$
$\quad+\underbrace{\frac{2\tilde{L}}{m}\sum_{i=1}^{m}\,\mathbb{E}\Bigg{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\Big{(}\nabla f_{i}(x_{t,k}^{i})-\nabla f_{i}(x_{t})\Big{)}\Bigg{\|}^{2}}_{\mathrm{A}_{4}}$
$\quad+\underbrace{\frac{2\tilde{L}}{m}\sum_{i=1}^{m}\,\mathbb{E}\Bigg{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\Big{(}\nabla f_{i}(x_{t})-\nabla f(x_{t})\Big{)}\Bigg{\|}^{2}}_{\mathrm{A}_{5}}$
$\quad+\underbrace{\frac{2\tilde{L}}{m}\sum_{i=1}^{m}\,\mathbb{E}\Bigg{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\nabla f(x_{t})\Bigg{\|}^{2}}_{\mathrm{A}_{6}}.$ (18)
Now, we provide an upper bound for each term in (18). Starting with $\mathrm{A}_{1},$ using (7), (8), and (11), we have
$\mathrm{A}_{1}$ $\leqslant\frac{K}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t,k}^{i})\Big{\|}^{2}$ (19)
$\leqslant\frac{K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}x_{t}-x_{t,k}^{i}\Big{\|}^{2}$
$=\frac{K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}$
$=\frac{K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\Big{(}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})+\nabla f_{i}(x_{t,\ell}^{i})\Big{)}\Big{\|}^{2}$
$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\Big{(}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{)}\Big{\|}^{2}+\,\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
$\leqslant\frac{2K^{2}\tilde{L}^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}}{D_{t}m}\sum_{i=1}^{m}\Big{(}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}^{2}.$ (20)
For $\mathrm{A}_{2},$ using (7) and (8), we have:
$\mathrm{A}_{2}$ $\leqslant\frac{K}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\nabla f_{i}(x_{t})-\nabla f(x_{t})\Big{\|}^{2}$
$\leqslant\frac{K\rho}{D_{t}m}\Big{\|}\nabla f(x_{t})\Big{\|}^{2}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}.$ (21)
For $\mathrm{A}_{3}$, using (7), (8), and (11), we have:
$\mathrm{A}_{3}$ $\leqslant\frac{2K\tilde{L}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,k}^{i})-\nabla f_{i}(x_{t,k}^{i})\Big{\|}^{2}$
$\leqslant\frac{2K\tilde{L}\sigma^{2}}{mb}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}.$ (22)
For $\mathrm{A}_{4},$ using (7), (8), and (11), we have:
$\mathrm{A}_{4}$ $\leqslant\frac{2K\tilde{L}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t,k}^{i})\Big{\|}^{2}$ (23)
$\leqslant\frac{2K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}x_{t}-x_{t,k}^{i}\Big{\|}^{2}$
$=\frac{2K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}$
$=\frac{2K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\Big{(}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})+\nabla f_{i}(x_{t,\ell}^{i})\Big{)}\Big{\|}^{2}$
$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\Big{(}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{)}\Big{\|}^{2}+\,\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
$\leqslant\frac{4K^{2}\tilde{L}^{3}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}}{m}\sum_{i=1}^{m}\Big{(}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}^{2}.$ (24)
For $\mathrm{A}_{5},$ using (7), (8), and (11), we have:
$\mathrm{A}_{5}$ $\leqslant\frac{2K\tilde{L}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\,\mathbb{E}\Big{\|}\nabla f_{i}(x_{t})-\nabla f(x_{t})\Big{\|}^{2}$
$\leqslant\frac{2K\tilde{L}\rho}{m}\Big{\|}\nabla f(x_{t})\Big{\|}^{2}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}.$ (25)
For $\mathrm{A}_{6},$ using (7), (8), and (11), we have:
$\mathrm{A}_{6}$ $\leqslant\frac{2K\tilde{L}}{m}\Big{\|}\nabla f(x_{t})\Big{\|}^{2}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}.$ (26)
Putting all bounds together.
Now, by putting (19)–(26) back into (18) and rearranging the terms, we obtain the following inequality:
$\Big{(}\frac{D_{t}}{2}-\frac{K\rho}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}-\frac{2K\tilde{L}(1+\rho)}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}\|\nabla f(x_{t})\|^{2}$
$\qquad\qquad\leqslant f(x_{t})-f(x_{t+1})+\frac{2K^{2}\tilde{L}^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}}{D_{t}m}\sum_{i=1}^{m}\Big{(}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}^{2}$
$\qquad\qquad\qquad+\frac{2K\tilde{L}\sigma^{2}}{mb}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}+\frac{4K^{2}\tilde{L}^{3}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}}{m}\sum_{i=1}^{m}\Big{(}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}^{2}.$ (27)
Now, it remains to notice that, by definition, $\eta_{t,k}^{i}=\mathcal{O}(\gamma)$, for all $i\in[m]$, $k\in\{0,1,\dots K{-}1\}$, and $t\geqslant 0$. Therefore, $D_{t}=\mathcal{O}(\gamma K)$. Thus, by dividing the above inequality by $\frac{D_{t}}{2}$, we have:
$\Big{(}1-\underbrace{\frac{2K\rho}{D_{t}^{2}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}}_{\mathcal{O}(1)}~{}-~{}\underbrace{\frac{4K\tilde{L}(1+\rho)}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}}_{\mathcal{O}(\gamma K)}\Big{)}\|\nabla f(x_{t})\|^{2}$
$\qquad\qquad\leqslant\underbrace{\frac{2(f(x_{t})-f(x_{t+1}))}{D_{t}}}_{\mathcal{O}\Big{(}\frac{1}{\gamma K}\Big{)}}~{}+~{}\underbrace{\frac{4K\tilde{L}\sigma^{2}}{D_{t}mb}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}}_{\mathcal{O}(\gamma K)}$
$\qquad\qquad\qquad+\underbrace{\frac{4K^{2}\tilde{L}^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}}{D_{t}^{2}m}\sum_{i=1}^{m}\Big{(}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}^{2}}_{\mathcal{O}(\gamma^{2}K^{2})}$
$\qquad\qquad\qquad+\underbrace{\frac{8K^{2}\tilde{L}^{3}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}}{D_{t}m}\sum_{i=1}^{m}\Big{(}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}^{2}}_{\mathcal{O}(\gamma^{3}K^{3})},$ (28)
for $\rho=\mathcal{O}(1)$. For the simplicity of exposition, assuming $\rho\leqslant\frac{1}{4}$ and $T\geqslant(5/\rho)^{2}=400$, we obtain the below by averaging (28) for $t=0,1,\dots,T{-}1$:
$\frac{1}{T}\sum_{t=0}^{T{-}1}\|\nabla f(x_{t})\|^{2}$ $\leqslant\mathcal{O}\Big{(}\frac{f(x_{0})-f^{\star}}{\gamma KT}\Big{)}+\mathcal{O}\Big{(}\frac{\gamma K\sigma^{2}}{b}\Big{)}+\mathcal{O}\Big{(}\tilde{L}^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}\gamma^{2}K^{2}\Big{)}$
$\quad+\mathcal{O}\Big{(}\tilde{L}^{3}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}\gamma^{3}K^{3}\Big{)}.$ (29)
Finally, by choosing $\gamma=\mathcal{O}\Big{(}\frac{1}{K\sqrt{T}}\Big{)}$, and defining $\Psi_{1}=\max\left\{\frac{\sigma^{2}}{b},f(x_{0})-f(x^{\star})\right\}$ and $\Psi_{2}=\left(\frac{\sigma^{2}}{b}+G^{2}\right)$ with $b=|\mathcal{B}|$ being the batch size, we arrive at the statement in Theorem 1, i.e.,
$\frac{1}{T}\sum_{t=0}^{T-1}\,\mathbb{E}\Big{\|}\nabla f\Big{(}x_{t}\Big{)}\Big{\|}^{2}$ $\leqslant\mathcal{O}\Big{(}\frac{\Psi_{1}}{\sqrt{T}}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{2}\Psi_{2}}{T}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{3}\Psi_{2}}{\sqrt{T^{3}}}\Big{)}.$
A.2 Proof of Algorithm 1 for convex case
Here, we provide an extension of the proof of (Malitsky and Mishchenko, 2020, Theorem 1) for distributed version. Note that the proof we provide here is not exactly the same version of $\Delta$-SGD presented in Algorithm 1; in particular, we assume no local iterations, i.e., $k=1$ for all $i\in[m]$, and every client participate, i.e, $p=1.$
We start with the extension of (Malitsky and Mishchenko, 2020, Lemma 1), which constructs a Lyapunov function for the distributed objective in (1).
Lemma 4.
Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex and twice differentiable. Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$, satisfy the following:
$\|x_{t+1}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\Big{[}\eta_{t}^{i}(1+\theta_{t}^{i})\big{(}f_{i}(x_{t}^{i})-f_{i}(x^{\star})\big{)}\Big{]}$
$\qquad\leqslant\|x_{t}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{t}^{i}-x_{t-1}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\Big{[}\eta_{t}^{i}\theta_{t}^{i}\big{(}f_{i}(x_{t-1}^{i})-f_{i}(x^{\star})\big{)}\Big{]}.$ (30)
The above lemma, which we prove below, constructs a contracting Lyapunov function, which can be seen from the second condition on the step size $\eta_{t}^{i}:$
$\eta_{t+1}^{i}\leqslant\sqrt{1+\theta_{t}^{i}}\eta_{t}^{i}\implies\eta_{t+1}^{i}\theta_{t+1}^{i}\leqslant(1+\theta_{t}^{i})\eta_{t}^{i}.$
Proof.
We denote $x_{t}=\frac{1}{m}\sum_{i=1}^{m}x_{t}^{i}.$
$\|x_{t+1}-x^{\star}\|^{2}$ $=\|x_{t}-\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\nabla f_{i}(x_{t}^{i})-x^{\star}\|$ (31)
$=\|x_{t}-x^{\star}\|^{2}+\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\nabla f_{i}(x_{t}^{i})\Big{\|}^{2}-2\Big{\langle}x_{t}-x^{\star},\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\nabla f_{i}(x_{t}^{i})\Big{\rangle}$ (32)
We will first bound the second term in (32). Using $\|\sum_{i=1}^{m}a_{i}\|^{2}\leqslant m\cdot\sum_{i=1}^{m}\|a_{i}\|^{2},$ we have
$\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\nabla f_{i}(x_{t}^{i})\Big{\|}^{2}$ $=\Big{\|}\frac{1}{m}\sum_{i=1}^{m}(x_{t}^{i}-x_{t+1}^{i})\Big{\|}^{2}\leqslant\frac{1}{m}\sum_{i=1}^{m}\|x_{t}^{i}-x_{t+1}^{i}\|^{2}.$ (33)
To bound $\|x_{t}^{i}-x_{t+1}^{i}\|^{2},$ observe that
$\|x_{t}^{i}-x_{t+1}^{i}\|^{2}$ $=2\|x_{t}^{i}-x_{t+1}^{i}\|^{2}-\|x_{t}^{i}-x_{t+1}^{i}\|^{2}$
$=2\langle-\eta_{t}^{i}\nabla f_{i}(x_{t}^{i}),x_{t}^{i}-x_{t+1}^{i}\rangle-\|x_{t}^{i}-x_{t+1}^{i}\|^{2}$
$=2\eta_{t}^{i}\langle\nabla f_{i}(x_{t}^{i})-\nabla f_{i}(x_{t-1}^{i}),x_{t+1}^{i}-x_{t}^{i}\rangle-\|x_{t}^{i}-x_{t+1}^{i}\|^{2}$
$\quad+2\eta_{t}^{i}\langle\nabla f_{i}(x_{t-1}^{i}),x_{t+1}^{i}-x_{t}^{i}\rangle$ (34)
For the first term in (34), using Cauchy-Schwarz and Young’s inequalities as well as
$\eta_{t}^{i}\leqslant\frac{\|x_{t}^{i}-x_{t-1}^{i}\|}{2\|\nabla f_{i}(x_{t}^{i})-\nabla f_{i}(x_{t-1}^{i})\|}$, we have:
$2\eta_{t}^{i}\langle\nabla f_{i}(x_{t}^{i})-\nabla f_{i}(x_{t-1}^{i}),x_{t+1}^{i}-x_{t}^{i}\rangle$ $\leqslant 2\eta_{t}^{i}\|\nabla f_{i}(x_{t}^{i})-\nabla f_{i}(x_{t-1}^{i})\|\cdot\|x_{t+1}^{i}-x_{t}^{i}\|$
$\leqslant\|x_{t}^{i}-x_{t-1}^{i}\|\cdot\|x_{t+1}^{i}-x_{t}^{i}\|$
$\leqslant\frac{1}{2}\|x_{t}^{i}-x_{t-1}^{i}\|^{2}+\frac{1}{2}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}.$
For the third term in (34), by convexity of $f_{i}(x),$ we have
$2\eta_{t}^{i}\langle\nabla f_{i}(x_{t-1}^{i}),x_{t+1}^{i}-x_{t}^{i}\rangle$ $=2\tfrac{\eta_{t}^{i}}{\eta_{t-1}^{i}}\langle x_{t-1}^{i}-x_{t}^{i},x_{t+1}^{i}-x_{t}^{i}\rangle$
$=2\theta_{t}^{i}\eta_{t}^{i}\langle x_{t-1}^{i}-x_{t}^{i},\nabla f_{i}(x_{t}^{i})\rangle$
$\leqslant 2\theta_{t}^{i}\eta_{t}^{i}[f_{i}(x_{t-1}^{i})-f_{i}(x_{t}^{i})].$
Together, we have
$\frac{1}{m}\sum_{i=1}^{m}\|x_{t}^{i}-x_{t+1}^{i}\|^{2}\leqslant\frac{1}{m}\sum_{i=1}^{m}\Big{\{}\tfrac{1}{2}\|x_{t}^{i}-x_{t-1}^{i}\|^{2}-\tfrac{1}{2}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}+2\theta_{t}^{i}\eta_{t}^{i}[f_{i}(x_{t-1}^{i})-f_{i}(x_{t}^{i})]\Big{\}}.$ (35)
Now to bound the last term in (32), we have
$-2\Big{\langle}x_{t}-x^{\star},\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\nabla f_{i}(x_{t}^{i})\Big{\rangle}$ $=\frac{2}{m}\eta_{t}^{i}\sum_{i=1}^{m}\Big{\langle}x^{\star}-x_{t}^{i},\nabla f_{i}(x_{t}^{i})\Big{\rangle}$
$\leqslant\frac{2}{m}\sum_{i=1}^{m}\eta_{t}^{i}\big{(}f_{i}(x^{\star})-f_{i}(x_{t}^{i})\big{)},$ (36)
where in the equality we used the fact that $x_{t}^{i}=x_{t},$ for $k=1.$
Putting (35) and (36) back to (32), we have
$\|x_{t+1}-x^{\star}\|^{2}$ $=\|x_{t}-x^{\star}\|^{2}+\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\nabla f_{i}(x_{t}^{i})\Big{\|}^{2}-2\Big{\langle}x_{t}-x^{\star},\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\nabla f_{i}(x_{t}^{i})\Big{\rangle}$
$\leqslant\|x_{t}-x^{\star}\|^{2}+\frac{1}{m}\sum_{i=1}^{m}\Big{\{}\tfrac{1}{2}\|x_{t}^{i}-x_{t-1}^{i}\|^{2}-\tfrac{1}{2}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}+2\theta_{t}^{i}\eta_{t}^{i}[f_{i}(x_{t-1}^{i})-f(x_{t}^{i})]\Big{\}}$
$\quad\quad-2\Big{\langle}x_{t}-x_{t}^{i},\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\nabla f_{i}(x_{t}^{i})\Big{\rangle}+2\frac{1}{m}\sum_{i=1}^{m}\eta_{t}^{i}\big{(}f_{i}(x^{\star})-f_{i}(x_{t}^{i})\big{)}.$ (37)
Averaging (37) for $i=1,\dots,m,$ notice that the first term in (37) disappears, as $\hat{x_{t}}=\frac{1}{m}\sum_{i=1}^{m}x_{i}.$
$\|x_{t+1}-x^{\star}\|^{2}+\frac{1}{m}\sum_{i=1}^{m}\Big{[}2\eta_{t}^{i}(1+\theta_{t}^{i})\big{(}f_{i}(x_{t}^{i})-f_{i}(x^{\star})\big{)}+\frac{1}{2}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}\Big{]}$
$\hskip 56.9055pt\leqslant\|x_{t}-x^{\star}\|^{2}+\frac{1}{m}\sum_{i=1}^{m}\Big{[}2\eta_{t}^{i}\theta_{t}^{i}\big{(}f_{i}(x_{t-1}^{i})-f_{i}(x^{\star})\big{)}+\frac{1}{2}\|x_{t}^{i}-x_{t-1}^{i}\|^{2}\Big{]}.$
Theorem 5.
Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex and twice differentiable. Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$, satisfy the following:
$\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}\leqslant\frac{2DL^{2}}{2Lt+1},$ (38)
where
$\tilde{x}_{t}^{i}$ $=\frac{\eta_{t}^{i}(1+\theta_{t}^{i})x_{t}^{i}+\sum_{k=1}^{t-1}\alpha_{k}^{i}x_{k}^{i}}{S_{t}^{i}},$
$\alpha_{k}^{i}$ $=\eta_{k}^{i}(1+\theta_{k}^{i})-\eta_{k+1}^{i}\theta_{k+1}^{i}$
$S_{t}^{i}$ $=\eta_{t}^{i}(1+\theta_{t}^{i})+\sum_{k=1}^{t-1}\alpha_{k}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}.$
Further, if $x^{\star}$ is any minimum of $f_{i}(x)$ for all $i\in[m]$, (38) implies convergence for problem (1).
Proof.
We start by telescoping (30). Then, we arrive at
$\|x_{t+1}-x^{\star}\|^{2}$ $+\frac{1}{2m}\sum_{i=1}^{m}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\Big{[}\eta_{t}^{i}(1+\theta_{t}^{i})\big{(}f_{i}(x_{t}^{i})-f_{i}(x^{\star})\big{)}$
$\hskip 99.58464pt+\sum_{k=1}^{t-1}\big{(}\eta_{k}^{i}(1+\theta_{k}^{i})-\eta_{k+1}^{i}\theta_{k+1}^{i}\big{)}\big{(}f_{i}(x_{k}^{i})-f_{i}(x^{\star})\big{)}\Big{]}$ (39)
$\leqslant\|x_{1}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{1}^{i}-x_{0}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\eta_{1}^{i}\theta_{1}^{i}\big{(}f_{i}(x_{0}^{i})-f_{i}(x^{\star})\big{)}:=D.$ (40)
Observe that (39) is of the form $\alpha_{t}^{i}[f_{i}(x_{t}^{i})-f_{i}(x^{\star})]+\sum_{k=1}^{t-1}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]=\sum_{k=1}^{t}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]$. The sum of the coefficients equals:
$\sum_{k=1}^{t}\alpha_{k}^{i}=\eta_{t}^{i}(1+\theta_{t}^{i})+\sum_{k=1}^{t-1}[\eta_{k}^{i}(1+\theta_{k}^{i})-\eta_{k+1}^{i}\theta_{k+1}^{i}]=\eta_{1}^{i}\theta_{1}^{i}+\sum_{k=1}^{t}\eta_{k}^{i}:=S_{t}^{i}.$ (41)
Therefore, by Jensen’s inequality, we have
$D\geqslant\|x_{1}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{1}^{i}-x_{0}^{i}\|^{2}+2S_{t}^{i}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}\geqslant 2S_{t}^{i}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)},$ (42)
which implies $f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\leqslant\frac{D^{i}}{2S_{t}^{i}}$, where
$\tilde{x}_{t}^{i}$ $=\frac{\eta_{t}^{i}(1+\theta_{t}^{i})x_{t}^{i}+\sum_{k=1}^{t-1}\alpha_{k}^{i}x_{k}^{i}}{S_{t}^{i}},$
$\alpha_{k}^{i}$ $=\eta_{k}^{i}(1+\theta_{k}^{i})-\eta_{k+1}^{i}\theta_{k+1}^{i}$
$S_{t}^{i}$ $=\eta_{t}^{i}(1+\theta_{t}^{i})+\sum_{k=1}^{t-1}\alpha_{k}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}.$ (43)
Now, notice that by definition of $\eta_{t}^{i},$ we have
$\eta_{t}^{i}=\frac{\|x_{t}^{i}-x_{t-1}^{i}\|}{2\|\nabla f(x_{t}^{i})-\nabla f(x_{t-1}^{i})\|}\geqslant\frac{1}{2\tilde{L}_{i}}\geqslant\frac{1}{2L},\quad\forall i\in[m].$
where $\tilde{L}_{i}$ is defined in (10), which by definition is smaller than or equal to the global smoothness $L$. Also, notice that $\eta_{1}^{i}\theta_{1}^{i}=\frac{(\eta_{1}^{i})^{2}}{\eta_{0}^{i}}=(\eta_{1}^{i})^{2},$ assuming for simplicity that $\eta_{0}^{i}=1.$ Thus, going back to (43), we have
$S_{t}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}$ $\geqslant\sum_{k=1}^{t}\frac{1}{2\tilde{L}_{i}}+\frac{1}{4\tilde{L}_{i}^{2}}$
$=\frac{t}{2L_{i}}+\frac{1}{4L_{i}^{2}}\geqslant\frac{t}{2L}+\frac{1}{4L^{2}}.$
Finally, we have
$\frac{D}{2}$ $\geqslant\frac{1}{m}\sum_{i=1}^{m}S_{t}^{i}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}$
$\geqslant\frac{1}{m}\sum_{i=1}^{m}\Big{(}\frac{t}{2L}+\frac{1}{4L^{2}}\Big{)}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)},$
which implies
$\frac{2DL^{2}}{2Lt+1}\geqslant\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}.$
A.3 Proof of adaptive gradient descent in nonconvex case: centralized
For the centralized version, we use the original step size from Malitsky and Mishchenko (2020) as in (4).
Theorem 6.
Running gradient descent with step size in (4) for nonconvex $f$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(x_{t})\|^{2}\leqslant\frac{8\tilde{L}}{3T}[f(x_{0})-f^{\star}],$
[…]
$\|\nabla f(x)-\nabla f(y)\|\leqslant\tilde{L}\|x-y\|\quad\forall x,y\in\mathcal{C}$
Proof.
Using Lemma 2, where with slight abuse of notation we interpret $f_{i}$ as $f$ and $\eta_{t}^{i}$ as $\eta_{t},$ we have
$f(x_{t+1})$ $\leqslant f(x_{t})+\langle\nabla f(x_{t}),x_{t+1}-x_{t}\rangle+\frac{1}{4\eta_{t}}\|x_{t+1}-x_{t}\|^{2}$
$=f(x_{t})+\langle\nabla f(x_{t}),-\eta_{t}\nabla f(x_{t})\rangle+\frac{1}{4\eta_{t}}\|x_{t+1}-x_{t}\|^{2}$
$=f(x_{t})-\eta_{t}\|\nabla f(x_{t})\|^{2}+\frac{\eta_{t}}{4}\|\nabla f(x_{t})\|^{2}$
$=f(x_{t})-\frac{3\eta_{t}}{4}\|\nabla f(x_{t})\|^{2}$
Rearranging, we have
$\|\nabla f(x_{t})\|^{2}\leqslant\frac{4}{3\eta_{t}}[f(x_{t})-f^{\star}+f^{\star}-f(x_{t-1})]$
Unfolding for $T$ iterations and using telescoping, we get
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(x_{t})\|^{2}\leqslant\frac{4}{3T\eta_{t}}[f(x_{0})-f^{\star}]\leqslant\frac{8\tilde{L}}{3T}[f(x_{0})-f^{\star}]$
as desired. ∎
Appendix B Additional Experiments and Plots
B.1 Ease of tuning
In this subsection, we show the effect of using different parameter $\delta$ in the step size of $\Delta$-SGD. As mentioned in Section 4, for $\Delta$-SGD, we append $\delta$ in front of the second condition, and hence the step size becomes
$\eta_{t,k}^{i}=\min\Big{\{}\tfrac{\gamma\|x_{t,k}^{i}-x_{t,{k-1}}^{i}\|}{2\|\tilde{\nabla}f_{i}(x_{t,k}^{i})-\tilde{\nabla}f_{i}(x_{t,{k-1}}^{i})\|},\sqrt{1+\delta\theta_{t,k-1}^{i}}\eta_{t,k-1}^{i}\Big{\}}.$
The experimental results presented in Table 1 were obtained using a value of $\delta=0.1$. For $\gamma$, we remind the readers that we did not change from the default value $\gamma=2$ in the original implementation in https://github.com/ymalitsky/adaptive_GD/blob/master/pytorch/optimizer.py.
In Figure 4, we display the final accuracy achieved when using different values of $\delta$ from the set $\{0.01,0.1,1\}$. Interestingly, we observe that $\delta$ has very little impact on the final accuracy. This suggests that $\Delta$-SGD is remarkably robust and does not require much tuning, while consistently achieving higher test accuracy compared to other methods in the majority of cases as shown in Table 1.
Figure 4: Effect of using different $\delta$ in the second condition of the step size of $\Delta$-SGD.
B.2 Illustration of the degree of non-iidness
As mentioned in the main text, the level of "non-iidness" is controlled using latent Dirichlet allocation (LDA) applied to the labels. This approach is based on Hsu et al. (2019) and involves assigning each client a multinomial distribution over the labels. This distribution determines the labels from which the client’s training examples are drawn. Specifically, the local training examples for each client are sampled from a categorical distribution over $N$ classes, parameterized by a vector $\mathbf{q}$. The vector $\mathbf{q}$ is drawn from a Dirichlet distribution with a concentration parameter $\alpha$ times a prior distribution vector $\mathbf{p}$ over $N$ classes.
To vary the level of "non-iidness," different values of $\alpha$ are considered: $0.01,0.1$, and $1$. A larger $\alpha$ corresponds to settings that are closer to i.i.d. scenarios, while smaller values introduce more diversity and non-iidness among the clients, as visualized in Figure 5. As can be seen, for $\alpha=1,$ each client (each row on $y$-axis) have fairly diverse classes, similarly to the i.i.d. case; as we decrease $\alpha,$ the number of classes that each client has access to decreases. With $\alpha=0.01,$ there are man y clients who only have access to a single or a couple classes.
Figure 5: Illustration of the degree of heterogeneity induced by using different concentration parameter $\alpha$ for Dirichlet distribution, for CIFAR-10 dataset (10 colors) and 100 clients (100 rows on $y$-axis).
B.3 Effect of different number of local iterations
In this subsection, we show the effect of different number of local epochs performed per client, while fixing the experimental setup to be the classification of CIFAR-100 datset, trained with a ResNet-50. We show for all three cases of the Dirichlet concentration parameter $\alpha\in\{0.01,0.1,1\}$. The result is shown in Figure 6.
As mentioned in the main text, instead of performing local iterations, we instead perform local epochs $E,$ similarly to Reddi et al. (2021); Li et al. (2020). All the results in Table 1 use $E=1,$ and in Figure 6, we show the results for $E\in\{1,2,3\}$. Note that in terms of $E$, the considered numbers might be small difference, but in terms of actual local gradient steps, they differ significantly.
As can be seen, using higher $E$ leads to slightly faster convergence in all cases of $\alpha.$ However, the speed up seems marginal compared to the amount of extra computation time it requires (i.e., using $E=2$ takes roughly twice more total wall-clock time than using $E=1$). Put differently, $\Delta$-SGD performs well with only using a single local epoch per client ($E=1$), and still can achieve great performance.
Figure 6:
B.4 Additional experiments using different number of local data per client
In this subsection, we provide additional experimental results, where the size of the local dataset each client has differs. In particular, instead of distributing $500$ samples to each client as done in Section 4 of the main text, we distribute $n_{i}\in[100,500]$ samples for each client $i$, where $n_{i}$ is the number of samples that client $i$ has, and is randomly sampled from the range $[100,500]$. We keep all other settings the same.
Then, in the server aggregation step (line 13 of Algorithm 1), we compute the weighted average of the local parameters, instead of the plain average, as suggested in McMahan et al. (2017), and used for instance in Reddi et al. (2021).
We experiment in two settings: CIFAR-10 dataset classification trained with a ResNet-18 He et al. (2016), with Dirichlet concentration parameter $\alpha=0.1,$ which was the setting where the algorithms were fined-tuned using grid search; please see details in Section 4. We also provide results for CIFAR-100 dataset classification trained with a ResNet-50, with the same $\alpha.$
The result can be found in Table 2. Similarly to the case with same number of data per client, $i.e.,n_{i}=500$ for all $i\in[m],$ we see that $\Delta$-SGD achieves good performance without any additional tuning. In particular, for CIFAR-10 classification trained with a ResNet-18, $\Delta$-SGD achieves the second best performance after SGD (without LR decay), with a small margin. Interestingly, SGDM (both with and without LR decay) perform much worse than SGD. This contrasts with the same setting in Table 1, where SGDM performed better than SGD, which again complicates how one should tune client optimizers in the FL setting. For CIFAR-100 classification trained with ResNet-50, $\Delta$-SGD outperforms all other methods with large margin.
Non-iidness Optimizer Dataset / Model
$\text{Dir}(\alpha\cdot\mathbf{p})$ CIFAR-10 CIFAR-100
ResNet-18 ResNet-50
$\alpha=1$ SGD 80.7${}_{\downarrow(0.0)}$ 53.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.0)}}$
SGD ($\downarrow$)↓ ) 78.8${}_{\downarrow(1.9)}$ 53.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.9)}}$
SGDM ${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.7)}}$ 53.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.6)}}$
SGDM ($\downarrow$)↓ ) 66.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.1)}}$ 53.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.4)}}$
Adam 79.9${}_{\downarrow(0.8)}$ 51.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.4)}}$
Adagrad 79.3${}_{\downarrow(1.4)}$ 44.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.0)}}$
SPS 64.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.3)}}$ 37.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(20.3)}}$
$\Delta$-SGD 80.4${}_{\downarrow(0.3)}$ 57.5${}_{\downarrow(0.0)}$
Table 2: Experimental results using different number of local data per client. Performance difference within 0.5% of the best result for each task are shown in bold. Subscripts${}_{\downarrow(x.x)}$ is the performance difference from the best result and is highlighted in pink when the difference is bigger than 2%. The symbol ($\downarrow$) appended to SGD and SGDM indicates step-wise learning rate decay, where the step sizes are divided by $10$ after 50%, and another by $10$ after 75% of the total rounds.

### v3

@@
 Federated learning (FL) is a distributed machine learning framework where the global model of a central server is trained via multiple collaborative steps by participating clients without sharing their data.
-While being a flexible framework, where the distribution of local data, participation rate, and computing power of each client can greatly vary, such flexibility gives rise to many new challenges, especially in the hyperparameter tuning on both the server and the client side.
+While being a flexible framework, where the distribution of local data, participation rate, and computing power of each client can greatly vary, such flexibility gives rise to many new challenges, especially in the hyperparameter tuning on the client side.
 We propose $\Delta$-SGD, a simple step size rule for SGD that enables each client to use its own step size by adapting to the local smoothness of the function each client is optimizing.
 We provide theoretical and empirical results where the benefit of the client adaptivity is shown in various FL scenarios.
-In particular, our proposed method achieves TOP-1 accuracy in 73% and TOP-2 accuracy in 100% of the experiments considered without additional tuning.
 1 Introduction
-Federated learning (FL) is a distributed machine learning framework that allows multiple clients to learn a global model collaboratively.
-Each client trains the model on its local data and sends only the updated model parameters to a central server for aggregation.
+Federated Learning (FL) is a machine learning framework that enables multiple clients to collaboratively learn a global model in a distributed manner.
+Each client trains the model on their local data, and then sends only the updated model parameters to a central server for aggregation.
 Mathematically, FL aims to solve the following optimization problem:
@@
 where $f_{i}(x):=\mathbb{E}_{z\sim\mathcal{D}_{i}}[F_{i}(x,z)]$ is the loss function of the $i$-th client, and $m$ is the number of clients.
-A key property of FL is its flexibility in how different clients participate in the overall training: the number of clients $m$, their participation rates, and computing power available to each client can all vary and change at any point during the overall training procedure.
-On top of that, each client’s local data is not shared with other clients or the server, resulting in certain inherent data privacy McMahan et al.
-(2017); Agarwal et al.
-(2018).
-While advantageous, such flexibility and data privacy also introduce a plethora of new challenges, notably: $i)$ how the server aggregates the local information coming from each client, and $ii)$ how to make sure each client meaningfully “learns” using their local data and computing device.
+A key property of FL its flexibility in how various clients participate in the overall training procedure.
+The number of clients, their participation rates, and computing power available to each client can vary and change at any time during the training.
+Additionally, the local data of each client is not shared with others, resulting in better data privacy (McMahan et al., 2017; Agarwal et al., 2018).
+While advantageous, such flexibility also introduces a plethora of new challenges, notably: $i)$ how the server aggregates the local information coming from each client, and $ii)$ how to make sure each client meaningfully “learns” using their local data and computing device.
 The first challenge was partially addressed in Reddi et al.
-(2021), where adaptive optimization methods, such as Adagrad Duchi et al.
-(2011) or Adam Kingma and Ba (2014), were utilized in the aggregation step.
-However, the second challenge remains largely unaddressed as we elaborate below, other than the exception of Wang et al.
-(2021); Xie et al.
-(2019) reviewed in Section 2.
-Local data of each client is not shared, which intrinsically implies heterogeneity in terms of the size and the distribution $\mathcal{D}_{i}$ of local datasets.
+(2021), where adaptive optimization methods such as Adam (Kingma & Ba, 2014) was utilized in the aggregation step.
+Yet, the second challenge remains largely unaddressed.
+Local data of each client is not shared, which intrinsically introduces heterogeneity in terms of the size and the distribution of local datasets.
 That is, $\mathcal{D}_{i}$ differs for each client $i$, as well as the number of samples $z\sim\mathcal{D}_{i}$.
-Consequently, $f_{i}(x)$ can vastly differ from client to client, making Problem (1) hard to solve (i.e., find an approximate minimizer).
-Moreover, the sheer amount of local updates is far larger than the number of aggregation steps, given that less communication is desired (i.e., less server-side aggregation) due to the high communication cost—typically 3-4$\times$ orders of magnitude more expensive than local computation—in distributed optimization settings Lan et al.
-(2020).
-Figure 1: (A): CIFAR-10 classification task trained on ResNet-18.
-(B): MNIST classification task trained on shallow CNN.
-(C): CIFAR-100 classification task trained on ResNet-50.
-All experiments use FedAvg McMahan et al.
-(2017) as the server optimizer.
-Extensive fine-tuning of the client optimizers is often required to achieve good performance.
-For instance, experimental results of the famous FedAvg algorithm were obtained after performing a grid-search of typically 11-13 step sizes of the clients’ SGD (McMahan et al., 2017, Section 3), as SGD (and its variants) are highly sensitive to the step size Moulines and Bach (2011); Toulis and Airoldi (2017); Assran and Rabbat (2020); Kim et al.
-(2022).
+Consequently, $f_{i}(x)$ can vastly differ from client to client, making the problem in (1) hard to optimize.
+Moreover, the sheer amount of local updates is far larger than the number of aggregation steps, due to the high communication cost—typically 3-4$\times$ orders of magnitude more expensive than local computation—in distributed settings (Lan et al., 2020).
+Figure 1: Illustration of the effect of not properly tuning the client step sizes.
+In (A), each client optimizer uses the best step size from grid-search.
+Then, the same step size from (A) is intentionally used in settings (B) and (C).
+Only $\Delta$-SGD works well across all settings without additional tuning.
+As a result, extensive fine-tuning of the client optimizers is often required to achieve good performance in FL scenarios.
+For instance, experimental results of the well-known FedAvg algorithm were obtained after performing a grid-search of typically 11-13 step sizes of the clients’ SGD (McMahan et al., 2017, Section 3), as SGD (and its variants) are highly sensitive to the step size (Toulis & Airoldi, 2017; Assran & Rabbat, 2020; Kim et al., 2022b).
 Similarly, in Reddi et al.
@@
 We plot the progress of test accuracies for different client optimizers, where, for all the other test cases, we intentionally use the same step size rules that were fine-tuned for the task in Figure 1(A).
-There, we train a ResNet-18 for CIFAR-10 dataset classification within an FL setting, where the best step sizes were used for all algorithms after grid-search; we defer the experimental details to Section 4.
+There, we train a ResNet-18 for CIFAR-10 dataset classification within an FL setting, where the best step sizes for each (client) optimizer was used after grid-search; we defer the experimental details to Section 4.
 Hence, all methods perform reasonably well, although $\Delta$-SGD, our proposed method, achieves noticeably better test accuracy when compared to non-adaptive SGD variants –e.g., a 5% difference in final classification accuracy from SGDM– and comparable final accuracy only with adaptive SGD variants, like Adam and Adagrad.
-In Figure 1(B), we train a shallow CNN for MNIST data classification using the same step size rules.
+In Figure 1(B), we train a shallow CNN for MNIST classification, using the same step sizes from (A).
 MNIST classification is now considered an “easy” task, and therefore SGD with the same constant and decaying step sizes from Figure 1(A) works well.
@@
 Adaptive optimizers, like Adam and Adagrad, show similar behavior, falling short in achieving good final accuracy, compared to their performance in the case of Figure 1(A).
-Similarly, in Figure 1(C), we plot the test accuracy progress for CIFAR-100 classification trained on ResNet-50, again using the same step size rules as before.
+Similarly, in Figure 1(C), we plot the test accuracy progress for CIFAR-100 classification trained on a ResNet-50, again using the same step size rules as before.
 Contrary to Figure 1(B), SGD with momentum (SGDM) works better than SGD, both with the constant and the decaying step sizes.
-Adam becomes a “good optimizer” again, but its “sibling”, Adagrad, performs worse than SGDM.
+Adam becomes a “good optimizer” again, but its “sibling,” Adagrad, performs worse than SGDM.
 On the contrary, our proposed method, $\Delta$-SGD, which we introduce in Section 3, achieves superior performance in all cases without any additional tuning.
@@
 Unfortunately, Figure 1 indicates that the answers to these questions highly vary depending on the setting; once the dataset itself or how the dataset is distributed among different clients changes, or once the model architecture changes, the client optimizers have to be properly re-tuned to ensure good performance.
-Perhaps surprisingly, the same holds for adaptive methods like Adagrad Duchi et al.
-(2011) and Adam Kingma and Ba (2014).
+Perhaps surprisingly, the same holds for adaptive methods like Adagrad (Duchi et al., 2011) and Adam (Kingma & Ba, 2014).
 Our hypothesis and contributions.
-Our paper takes a stab in this direction: we propose DELTA-SGD (DistributEd LocaliTy Adaptive SGD), a simple adaptive SGD scheme, that does not require any knowledge of the problem constants, such as the $L$-smoothness parameter.
+Our paper takes a stab in this direction: we propose DELTA-SGD (DistributEd LocaliTy Adaptive SGD), a simple adaptive distributed SGD scheme, that can automatically tune its step size based on the available local data.
 We will refer to our algorithm as $\Delta$-SGD in the rest of the text.
-Our adaptive step size is extended from Malitsky and Mishchenko (2020), where the focus was centralized convex optimization; we defer the readers to Section 3 for details.
 Our contributions can be summarized as follows:
-  • We extend the (centralized) adaptive step size from Malitsky and Mishchenko (2020) to the FL setting.
-The implication of this adaptive step size in FL is twofold: $i)$ each client can use its own step size, and $ii)$ each client’s step size adapts to the local smoothness of $f_{i}$—hence LocaliTy Adaptive— and can even increase during local iterations.
-Moreover, thanks to the simplicity of the proposed step size, our method is agnostic to the server optimizer as well as the loss function, and thus can be easily combined with server adaptive methods such as FedAdam Reddi et al.
-(2021), or methods that use different (regularized) loss functions such as FedProx Li et al.
-(2020) or MOON Li et al.
-(2021).
+  • We propose $\Delta$-SGD, which has two implications in FL settings: $i)$ each client can use its own step size, and $ii)$ each client’s step size adapts to the local smoothness of $f_{i}$ –hence LocaliTy Adaptive– and can even increase during local iterations.
+Moreover, due to the simplicity of the proposed step size, $\Delta$-SGD is agnostic to the loss function and the server optimizer, and thus can be combined with methods that use different loss functions such as FedProx (Li et al., 2020) or MOON (Li et al., 2021), or adaptive server methods such as FedAdam (Reddi et al., 2021).
   • We provide convergence analysis of $\Delta$-SGD in a general nonconvex setting (Theorem 1).
-We also provide the convergence analysis in the convex setting, extending the analysis from Malitsky and Mishchenko (2020) to a distributed one, as well as a simple centralized nonconvex convergence analysis, which was an open question posed in Malitsky and Mishchenko (2020).
-Due to space constraints, we only state the result of general nonconvex case in Theorem 1, and defer other theorems and proofs to Appendix A.
+We also prove convergence in convex setting; due to space constraints, we only state the result of the general nonconvex case in Theorem 1, and defer other theorems and proofs to Appendix A.
   • We evaluate our approach on several benchmark datasets and demonstrate that $\Delta$-SGD achieves superior performance compared to other state-of-the-art FL methods.
-Our experiments show that our method can effectively adapt the client step size to the underlying data distribution and achieve faster convergence of the global model, without any additional tuning.
+Our experiments show that $\Delta$-SGD can effectively adapt the client step size to the underlying local data distribution, and achieve convergence of the global model without any additional tuning.
 Our approach can help overcome the client step size tuning challenge in FL and enable more efficient and effective collaborative learning in distributed systems.
@@
 A bit of background on optimization theory.
-Arguably the most fundamental optimization algorithm for minimizing a function $f(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the gradient descent (GD) Cauchy et al.
-(1847), which iterates with the step size $\eta_{t}$ as: $x_{t+1}=x_{t}-\eta_{t}\nabla f(x_{t})$.
+Arguably the most fundamental optimization algorithm for minimizing a function $f(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the gradient descent (GD), which iterates with the step size $\eta_{t}$ as: $x_{t+1}=x_{t}-\eta_{t}\nabla f(x_{t})$.
 Under the assumption that $f$ is globally $L$-smooth, that is:
-$\|\nabla f(x)-\nabla f(y)\|\leqslant L\cdot\|x-y\|\quad\forall x,y,$ (2)
-the step size $\eta_{t}=\frac{1}{L}$ is the “optimal” step size for GD, which converges at the rate, which is not improvable Nesterov (2003); Drori and Teboulle (2014), for convex $f$:
-$f(x_{t+1})-f(x^{\star})\leqslant\frac{L\|x_{0}-x^{\star}\|^{2}}{2(2t+1)},$ (3)
-while $\eta_{t}\in(0,\frac{2}{L})$ defines the “admissible” step size range required to ensure that GD converges for this function class Polyak (1963).
-Similarly, if $f$ is also (globally) $\mu$-strongly convex, that is:
-$f(x)\geqslant f(y)+\langle\nabla f(y),x-y\rangle+\tfrac{\mu}{2}\|x-y\|^{2}~{}~{}\forall x,y,$
-then the optimal step size is $\eta_{t}=\frac{2}{\mu+L}$ Nesterov (2003), and the rate improves to a linear one:
-$\|x_{t+1}-x^{\star}\|^{2}\leqslant\Big{(}\frac{L-\mu}{L+\mu}\Big{)}^{t}\|x_{0}-x^{\star}\|^{2}.$
+$\|\nabla f(x)-\nabla f(y)\|\leqslant L\cdot\|x-y\|\quad\forall x,y\in\mathbb{R}^{d},$ (2)
+the step size $\eta_{t}=\frac{1}{L}$ is the “optimal” step size for GD, which converges for convex $f$ at the rate:
+$f(x_{t+1})-f(x^{\star})\leqslant\frac{L\|x_{0}-x^{\star}\|^{2}}{2(2t+1)}.$ (3)
 Limitations of popular step sizes.
-In practice, however, the constants $L$ and $\mu$ are rarely known Boyd et al.
-(2004).
-As such, there has been a plethora of efforts in the optimization community to come up with a general step size rule or method that does not require a priori knowledge of problem constants Nesterov (2015); Orabona and Pál (2016); Levy et al.
-(2018), so that an optimization algorithm can work as “plug-and-play.”
-This is a well-known research question, with some major lines of work on this direction being: $i)$ line-search Armijo (1966); Paquette and Scheinberg (2020), $ii)$ adaptive learning rate (e.g., Adagrad Duchi et al.
-(2011) and Adam Kingma and Ba (2014)), and $iii)$ Polyak step size Polyak (1969); Hazan and Kakade (2019), to name a few.
-Unfortunately, the search for the “perfect” step size is still active Loizou et al.
-(2021); Defazio et al.
-(2022); Defazio and Mishchenko (2023); Bernstein et al.
-(2023); Latafat et al.
-(2023), as the aforementioned approaches have limitations.
-For instance, to use line search, solving a sub-routine is often required, inevitably incurring additional evaluations of the function and/or the gradient.
-To use methods like Adagrad and Adam, the knowledge of the initial distance from the solution set $D$, which is nontrivial to find, is required to ensure good performance Defazio and Mishchenko (2023); see also Section 4.
-Lastly, Polyak step size requires a sharp knowledge of $f(x^{\star})$ Hazan and Kakade (2019), which is not always known a priori, and is unclear what value to use for an arbitrary loss function and model.
-Implications in FL scenarios.
-On top of that, even with the knowledge of $L$, GD can be slow.
-Indeed, the global $L$-smoothness condition, as described above, needs to be satisfied for all $x$ and $y$, implying even finite $L$ can be arbitrarily large.
+In practice, however, the constants like $L$ are rarely known (Boyd et al., 2004).
+As such, there has been a plethora of efforts in the optimization community to develop a universal step size rule or method that does not require a priori knowledge of problem constants so that an optimization algorithm can work as “plug-and-play” (Nesterov, 2015; Orabona & Pál, 2016; Levy et al., 2018).
+A major lines of work on this direction include: $i)$ line-search (Armijo, 1966; Paquette & Scheinberg, 2020), $ii)$ adaptive learning rate (e.g., Adagrad (Duchi et al., 2011) and Adam (Kingma & Ba, 2014)), and $iii)$ Polyak step size (Polyak, 1969), to name a few.
+Unfortunately, the pursuit of finding the “ideal” step size is still active (Loizou et al., 2021; Defazio & Mishchenko, 2023; Bernstein et al., 2023), as the aforementioned approaches have limitations.
+For instance, to use line search, solving a sub-routine is required, inevitably incurring additional evaluations of the function and/or the gradient.
+To use methods like Adagrad and Adam, the knowledge of $D$, the distance from the initial point to the solution set, is required to ensure good performance (Defazio & Mishchenko, 2023).
+Similarly, Polyak step size requires knowledge of $f(x^{\star})$ (Hazan & Kakade, 2019), which is not always known a priori, and is unclear what value to use for an arbitrary loss function and model; see also Section 4 where not properly estimating $D$ or $f(x^{\star})$ resulting in suboptimal performance of relevant client optimizers in FL scenarios.
+Implications in distributed/FL scenarios.
+The global $L$-smoothness in (2) needs to be satisfied for all $x$ and $y$, implying even finite $L$ can be arbitrarily large.
 This results in a small step size, leading to slow convergence, as seen in (3).
+Malitsky & Mishchenko (2020) attempted to resolve this issue by proposing a step size for (centralized) GD that depends on the local smoothness of $f$, which by definition is smaller than the global smoothness constant $L$ in (2).
+$\Delta$-SGD is inspired from this work.
 Naturally, the challenge is aggravated in the distributed case.
-To solve (1), under the assumption that each $f_{i}$ is $L_{i}$-smooth, the step size of the form ${1}/{L_{\text{max}}}$ where ${L_{\text{max}}}:=\max_{i}L_{i}$ is often used Yuan et al.
-(2016); Scaman et al.
-(2017); Qu and Li (2019), to ensure convergence Uribe et al.
-(2020).
-Yet, one can easily imagine a situation where $L_{i}\gg L_{j}$ for $i\neq j,$ in which case the convergence of $f_{j}(x)$ can be arbitrarily slow by using step size ${1}/{L_{\text{max}}}$.
-There are plenty of practical examples where such a scenario could happen.
-For instance, imagine smartphone photo albums of the members of a fishing club versus a bird-watching club.
-The albums of these two groups could have vastly different data distributions, one skewed with pictures including fishes and rivers, while the other skewed with birds and forests Wah et al.
-(2011).
-On top of that, smartphones from different manufacturers nowadays have distinctly different camera filters applied—recently a big selling point for companies like Apple, introducing technologies like “portrait mode,” and Samsung, introducing 100$\times$ zoom capability, both based on image processing techniques—making even similar pictures to have distinct pixel values, when taken with different smartphones.
-Therefore, to ensure that each client learns useful information in the FL setting as in (1), ideally: $i)$ each agent should be able to use its own step size, instead of crude ones like $1/L_{\text{max}}$ for all; $ii)$ the individual step size should be “locally adaptive” to the function $f_{i},$ (i.e., even $1/L_{i}$ can be too crude, analogously to the centralized GD); and $iii)$ the step size should not depend on problem constants like $L_{i}$ an $\mu_{i}$, which are often unknown.
+To solve (1) under the assumption that each $f_{i}$ is $L_{i}$-smooth, the step size of the form ${1}/{L_{\text{max}}}$ where ${L_{\text{max}}}:=\max_{i}L_{i}$ is often used (Yuan et al., 2016; Scaman et al., 2017; Qu & Li, 2019), to ensure convergence (Uribe et al., 2020).
+Yet, one can easily imagine a situation where $L_{i}\gg L_{j}$ for $i\neq j,$ in which case the convergence of $f_{j}(x)$ can be arbitrarily slow by using step size of the form ${1}/{L_{\text{max}}}$.
+Therefore, to ensure that each client learns useful information in FL settings as in (1), ideally: $i)$ each agent should be able to use its own step size, instead of crude ones like $1/L_{\text{max}}$ for all agents; $ii)$ the individual step size should be “locally adaptive” to the function $f_{i},$ (i.e., even $1/L_{i}$ can be too crude, analogously to the centralized GD); and $iii)$ the step size should not depend on problem constants like $L_{i}$ or $\mu_{i}$, which are often unknown.
 In Section 3, we introduce DistributEd LocaliTy Adaptive SGD ($\Delta$-SGD), which satisfies all of the aforementioned desiderata.
 Related work on FL.
-FedAvg McMahan et al.
-(2017) is the simplest and probably the most famous FL algorithm, which takes a (weighted) average of the client parameters in the aggregation step after clients perform some local iterations.
-While FedAvg has achieved great results empirically, it has been observed recently that it can suffer from model drift Praneeth Karimireddy et al.
-(2019) induced by the heterogeneity Hsu et al.
+The FL literature is vast; here, we focus on the results that closely relate to our work.
+The FedAvg algorithm (McMahan et al., 2017) is one of the simplest FL algorithms, which averages client parameters after some local iterations.
+Reddi et al.
+(2021) showed that FedAvg is a special case of a meta-algorithm, where both the clients and the server use SGD optimizer, with the server learning rate being 1.
+To handle data heterogeneity and model drift, they proposed using adaptive optimizers at the server.
+This results in algorithms such as  FedAdam and FedYogi, which respectively use the Adam (Kingma & Ba, 2014) and Yogi (Zaheer et al., 2018) as server optimizers.
+Our approach is orthogonal and complimentary to this, as we propose a client adaptive optimizer that can be easily combined with these server-side aggregation methods (c.f., Appendix B.4).
+Other approaches have handled the heterogeneity by changing the loss function.
+FedProx (Li et al., 2020) adds an $\ell_{2}$-norm proximal term to to handle data heterogeneity.
+Similarly, MOON (Li et al., 2021) uses the model-contrastive loss between the current and previous models.
+Again, our proposed method can be seamlessly combined with these approaches (c.f., Table 4(b), Appendix B.6 and B.5).
+The closest related work to ours is a concurrent work by Mukherjee et al.
 (2023).
-At the core, simple averaging is too crude of an operation.
-As such, FedAvg uses a weighted average heuristic, where the weights are based on the number of data $n_{i}$ for each client $i$ McMahan et al.
+There, authors utilize the Stochastic Polyak Stepsize (Loizou et al., 2021) in FL settings.
+We do include SPS results in Section 4.
+There are some previous works that considered client adaptivity, including AdaAlter (Xie et al., 2019a) and Wang et al.
 (2021).
-The authors in Reddi et al.
-(2021) took this idea one step further: it showed that FedAvg can be seen as a special case where both the clients and the server use the SGD optimizer, with a server learning rate of 1.
-To handle the heterogeneity and the model drift, they proposed to use adaptive optimization scheme at the server, resulting in, e.g., FedAdam or FedYogi, which respectively use Adam Kingma and Ba (2014) and Yogi Zaheer et al.
-(2018) as the server optimizer, and can improve upon FedAvg in various scenarios.
-Note that this fundamentally differs from our approach, where we propose a client adaptive optimizer that can be easily combined with sophisticated server-side aggregation methods like FedAdam.
-Perhaps the closest works in spirit to our work are AdaAlter Xie et al.
-(2019) and locality adaptivity in FL Wang et al.
-(2021).
-Both works utilize an adaptive client optimizer, similar to Adagrad Duchi et al.
-(2011), with either delayed or restarted updates on the adaptive variables (e.g., pre-conditioners and momentum buffers).
-However, as noted in Reddi et al.
-(2021), AdaAlter incurs twice the memory compared to FedAvg, as AdaAlter requires communicating both the model parameters as well as the “client accumulators,” which is of the same dimension as model parameters.
-Wang et al.
-(2021) overcomes this issue by simply restarting the client optimizer states, but requires complicated local and global correction step to achieve good performance.
-On the contrary, $\Delta$-SGD changes virtually nothing compared to FedAvg other than the simple step size computation, and yet still achieves impressive performance; see Section 4.
-Other approaches have handled the heterogeneity by changing the loss function.
-FedProx Li et al.
-(2020) adds an $\ell_{2}$-norm proximal term to the individual loss functions such that the local parameters do not diverge too much from the previously communicated global model.
-Similarly, MOON Li et al.
-(2021) uses the cross-entropy loss between the current and previous models.
-As mentioned in the introduction, our proposed method can be seamlessly combined with these approaches.
-Another line of works attempts to handle heterogeneity by utilizing control-variates Praneeth Karimireddy et al.
-(2019); Liang et al.
-(2019); Karimireddy et al.
-(2020), a similar idea to variance reduction from single-objective optimization Johnson and Zhang (2013); Gower et al.
-(2020).
-While theoretically appealing, these methods either require full gradient computation periodically or are not usable when not all the clients participate, as the clients must maintain a state throughout all rounds.
+Both works utilize an adaptive client optimizer similar to Adagrad.
+However, AdaAlter incurs twice the memory compared to FedAvg, as AdaAlter requires communicating both the model parameters as well as the “client accumulators”; similarly, Wang et al.
+(2021) requires complicated local and global correction steps to achieve good performance.
+In short, previous works that utilize client adaptivity require modification in server-side aggregation; without such heuristics, Adagrad (Duchi et al., 2011) exhibits suboptimal performance, as seen in Table 1.
+Lastly, another line of works attempts to handle heterogeneity by utilizing control-variates (Praneeth Karimireddy et al., 2019; Liang et al., 2019; Karimireddy et al., 2020), a similar idea to variance reduction from single-objective optimization (Johnson & Zhang, 2013; Gower et al., 2020).
+While theoretically appealing, these methods either require periodic full gradient computation or are not usable under partial client participation, as all the clients must maintain a state throughout all rounds.
 3 DELTA($\Delta$)-SGD: Distributed locality adaptive SGD
-In Malitsky and Mishchenko (2020), a novel step size rule for (centralized) GD was introduced, which adapts to the minimization function’s local geometry.
-The proposed step size is clever and strikingly simple:
-$\eta_{t}=\min\Big{\{}\tfrac{\|x_{t}-x_{t-1}\|}{2\|\nabla f(x_{t})-\nabla f(x_{t-1})\|},\sqrt{1+\theta_{t-1}}\eta_{t-1}\Big{\}},\quad\theta_{t-1}=\eta_{t-1}/\eta_{t-2}.$ (4)
-The step size in (4) adapts to the local smoothness of the iterates, that is:
-$\|\nabla f(x_{t})-\nabla f(x_{t-1})\|\leqslant L_{t}\cdot\|x_{t}-x_{t-1}\|,\quad\forall t=1,2,\dots$ (5)
-and hence can increase during iterations, while the second condition in (4) ensures that $\eta_{t}$ does not increase arbitrarily.
-By definition, $L_{t}$ in (5) is never bigger than the global $L$ in (2), resulting in faster convergence in practice (Malitsky and Mishchenko, 2020, Section 4).
-Note that (4) is vastly different from some standard adaptive methods like Adagrad Duchi et al.
-(2011), which can only decrease the learning rate, as they accumulate non-negative quantities in the denominator of the step size.
-As a result, the scalar multiplied to their “adaptive” learning rates play a crucial role and requires proper tuning, similarly to vanilla SGD, as evident in Figure 1; see also Table 1 in Section 4.
-We now introduce DELTA($\Delta$)-SGD: DistributEd LocaliTy Adaptive SGD, where we extend (4) to the FL setting, with the inclusion of stochasticity and local iterations, as summarized in Algorithm 1.
-For brevity, we use ~$\tilde{\nabla}f_{i}(x)=\frac{1}{|\mathcal{B}|}\sum\limits_{z\in\mathcal{B}}F_{i}(x,z)$ to denote the stochastic gradients with batch size $|\mathcal{B}|=b.$
+We now introduce $\Delta$-SGD. In its simplest form, client $i$ at communication round $t$ uses the step size:
+$\eta_{t}^{i}=\min\Big{\{}\tfrac{\|x_{t}^{i}-x_{t-1}^{i}\|}{2\|\nabla f_{i}(x_{t}^{i})-\nabla f_{i}(x_{t-1}^{i})\|},\sqrt{1+\theta_{t-1}^{i}}\eta_{t-1}^{i}\Big{\}},\quad\theta_{t-1}^{i}=\eta_{t-1}^{i}/\eta_{t-2}^{i}.$ (4)
+The first part of $\min\{\cdot,\cdot\}$ approximates the (inverse of) local smoothness of $f_{i}$, [^1] Notice that $\|\nabla f_{i}(x_{t}^{i})-\nabla f_{i}(x_{t-1}^{i})\|\leqslant\tfrac{1}{2\eta_{t}^{i}}\|x_{t}^{i}-x_{t-1}^{i}\|\approx\tilde{L}_{i,t}\|x_{t}^{i}-x_{t-1}^{i}\|.$ and the second part controls how fast $\eta_{t}^{i}$ can increase.
+Indeed, $\Delta$-SGD with (4) enjoys the following decrease in the Lyapunov function:
+$\|x_{t+1}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\Big{[}\eta_{t+1}^{i}\theta_{t+1}^{i}\big{(}f_{i}(x_{t}^{i})-f_{i}(x^{\star})\big{)}\Big{]}$
+$\qquad\leqslant\|x_{t}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{t}^{i}-x_{t-1}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\Big{[}\eta_{t}^{i}\theta_{t}^{i}\big{(}f_{i}(x_{t-1}^{i})-f_{i}(x^{\star})\big{)}\Big{]},$ (5)
+when $f_{i}$’s are assumed to be convex (c.f., Theorem 5 in Appendix A).
+For the FL settings, we extend (4) by including stochasticity and local iterations, as summarized in Algorithm 1 and visualized in Figure 7.
+For brevity, we use ~$\tilde{\nabla}f_{i}(x)=\frac{1}{|\mathcal{B}|}\sum_{z\in\mathcal{B}}\nabla F_{i}(x,z)$ to denote the stochastic gradients with batch size $|\mathcal{B}|=b$.
 Algorithm 1 DELTA($\Delta$)-SGD: DistributEd LocaliTy Adaptive SGD
@@
 9:               $\eta_{t,k}^{i}=\min\Big{\{}\frac{\gamma\|x_{t,k}^{i}-x_{t,{k-1}}^{i}\|}{2\|\tilde{\nabla}f_{i}(x_{t,k}^{i})-\tilde{\nabla}f_{i}(x_{t,{k-1}}^{i})\|},\sqrt{1+\theta_{t,k-1}^{i}}\eta_{t,k-1}^{i}\Big{\}}$
-10:              $\theta_{t,k}^{i}=\eta_{t,k}^{i}/\eta_{t,k-1}^{i}$triangleright (line 10 & 11) update locality adaptive step size
+10:              $\theta_{t,k}^{i}=\eta_{t,k}^{i}/\eta_{t,k-1}^{i}$triangleright (line 9 & 10) update locality adaptive step size
 11:         end for
 12:     end for
-13:     $x_{t+1}=\frac{1}{|\mathcal{S}_{t}|}\sum\limits_{i\in\mathcal{S}_{t}}x_{t,K}^{i}$triangleright server-side aggregation
+13:     $x_{t+1}=\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}x_{t,K}^{i}$triangleright server-side aggregation
 14:end for
 15:return $x_{T}$
-We make a few remarks about Algorithm 1.
-First, the input $\eta_{0}>0$ can be quite arbitrary, as it can be corrected, per client level, in the 1st local iteration (line 9); similarly for $\theta_{0}>0$ (line 10).
-Second, as the step size $\eta_{t,k}^{i}$ requires two successive stochastic gradients, there could be two options: using $i)$ the same batch or $ii)$ a new batch; the second option will incur additional memory requirement, so we use the first.
-[^1] This distinction was already analyzed in the centralized setting, and it was reported that using the same batch performed better in practice Malitsky and Mishchenko (2020).
-Third, we include the “amplifier” $\gamma$ to the first condition of step size (line 9), but this is only needed for Theorem 1.222In practice, we did not change from the default value $\gamma=2$ in the original implementation in https://github.com/ymalitsky/adaptive_GD/blob/master/pytorch/optimizer.py.
-fourth, after local steps are done, $\eta_{0}$ is re-used (line 6); we could potentially use a more sophisticated approach and use a similar adaptive step size to line 9, but we leave this for future work.
-Last, we present simple averaging as the server-side aggregation for simplicity (line 13), but more sophisticated methods like FedAdam Reddi et al.
-(2021) can be used instead.
+We make a few remarks of Algorithm 1.
+First, the input $\theta_{0}>0$ can be quite arbitrary, as it can be corrected, per client level, in the first local iteration (line 10); similarly for $\eta_{0}>0$, although $\eta_{0}$ should be sufficiently small to prevent divergence in the first local step.
+Second, we include the “amplifier” $\gamma$ to the first condition of step size (line 9), but this is only needed for Theorem 1.222For all our experiments, we use the default value $\gamma=2$ from the original implementation in https://github.com/ymalitsky/adaptive_GD/blob/master/pytorch/optimizer.py.
+Last, ~$\tilde{\nabla}f_{i}(x_{t,k-1}^{i})$ shows up twice: in updating $x_{t,k}^{i}$ (line 8) and $\eta_{t,k}^{i}$ (line 9).
+Thus, one can use the same or different batches; we use the same batches in experiments to prevent additional gradient evaluations.
 3.1 Convergence Analysis
 Technical challenges.
-Before we present the convergence analysis, let us elaborate on the non-trivial technical challenges of analyzing Algorithm 1.
-The main difference of analyzing Algorithm 1 compared to other distributed/decentralized optimization theory is that the step size $\eta_{t,k}^{i}$ not only depends on the round $t$ and local steps $k$ (which is a challenge shared by other analyses), but also on $i$, the client.
-To deal with the client-dependent step size, we require a slightly non-standard assumption on the dissimilarity between $f$ and $f_{i}$, as shown below.
+The main difference of analyzing Algorithm 1 compared to other decentralized optimization algorithms is that the step size $\eta_{t,k}^{i}$ not only depends on the round $t$ and local steps $k$, but also on $i$, the client.
+To deal with the client-dependent step size, we require a slightly non-standard assumption on the dissimilarity between $f$ and $f_{i}$, as detailed below.
 Assumption 1.
 There exist nonnegative constants $\sigma,\rho,$ and $G$ such that for all $i\in[M]$ and $x\in\mathbb{R}^{d}$,
-$\mathbb{E}\|\nabla F_{i}(x,z)-\nabla f_{i}(x)\|$ $\leqslant\sigma^{2},$ (bounded variance) (1a)
+$\mathbb{E}\|\nabla F_{i}(x,z)-\nabla f_{i}(x)\|^{2}$ $\leqslant\sigma^{2},$ (bounded variance) (1a)
 $\|\nabla f_{i}(x)\|$ $\leqslant G,$ (bounded gradient) (1b)
 $\|\nabla f_{i}(x)-\nabla f(x)\|^{2}$ $\leqslant\rho\|\nabla f(x)\|^{2}.$ (strong growth of dissimilarity) (1c)
-Assumption 1a has been standard in stochastic optimization literature Moulines and Bach (2011); Ghadimi and Lan (2013); Stich (2018); Khaled et al.
-(2020); recently, this assumption has been relaxed to weaker ones such as weak growth assumption Bottou et al.
-(2018) or expected smoothness Gower et al.
-(2021), but this is out of scope of this work.
-Assumption 1b is fairly standard in nonconvex optimization Zaheer et al.
-(2018); Ward et al.
-(2020), and often used in FL setting Xie et al.
-(2019, 2019); Reddi et al.
-(2021).
-Assumption 1c is reminiscent of the strong growth assumption in stochastic optimization literature Schmidt and Roux (2013), which is still used in recent works Cevher and Vũ (2019); Vaswani et al.
-(2019); Vaswani et al.
-(2019).
-To the best of our knowledge, this is the first theoretical analysis of the FL setting where clients can use their step size.
-We now present the main convergence theorem of Algorithm 1.
+Assumption 1a has been standard in stochastic optimization literature (Ghadimi & Lan, 2013; Stich, 2018; Khaled et al., 2020); recently, this assumption has been relaxed to weaker ones such as expected smoothness (Gower et al., 2021), but this is out of scope of this work.
+Assumption 1b is fairly standard in nonconvex optimization (Zaheer et al., 2018; Ward et al., 2020), and often used in FL setting (Xie et al., 2019a; b; Reddi et al., 2021).
+Assumption 1c is reminiscent of the strong growth assumption in stochastic optimization (Schmidt & Roux, 2013), which is still used in recent works (Cevher & Vũ, 2019; Vaswani et al., 2019a; b).
+To the best of our knowledge, this is the first theoretical analysis of the FL setting where clients can use their own step size.
+We now present the main theorem.
 Theorem 1.
@@
 $\frac{1}{T}\sum_{t=0}^{T-1}\,\mathbb{E}\left\|\nabla f\left(x_{t}\right)\right\|^{2}$ $\leqslant\mathcal{O}\left(\frac{\Psi_{1}}{\sqrt{T}}\right)+\mathcal{O}\left(\frac{\tilde{L}^{2}\Psi_{2}}{T}\right)+\mathcal{O}\left(\frac{\tilde{L}^{3}\Psi_{2}}{\sqrt{T^{3}}}\right),$
-where $\Psi_{1}=\max\left\{\frac{\sigma^{2}}{b},f(x_{0})-f(x^{\star})\right\}$ and $\Psi_{2}=\left(\frac{\sigma^{2}}{b}+G^{2}\right)$ are global constants, with $b=|\mathcal{B}|$ being the batch size; $\tilde{L}$ is a constant at most the average of local smoothness, i.e., $\max_{t}\frac{1}{m}\sum_{i=1}^{m}L_{i,t},$ where $L_{i,t}$ the local smoothness of $f_{i}$ at round $t$.
+where $\Psi_{1}=\max\left\{\frac{\sigma^{2}}{b},f(x_{0})-f(x^{\star})\right\}$ and $\Psi_{2}=\left(\frac{\sigma^{2}}{b}+G^{2}\right)$ are global constants, with $b=|\mathcal{B}|$ being the batch size; $\tilde{L}$ is a constant at most the maximum of local smoothness, i.e., $\max_{i,t}\tilde{L}_{i,t},$ where $\tilde{L}_{i,t}$ the local smoothness of $f_{i}$ at round $t$.
 The convergence result in Theorem 1 implies a sublinear convergence to an $\varepsilon$-first order stationary point with at least $T=\mathcal{O}(\varepsilon^{-2})$ communication rounds.
-We emphasize again that the conditions on $\gamma$ and $\eta_{0}$ are only required for our theory.
-[^3] In all experiments in Section 4, we use the default setting $\gamma=2$, and $\eta_{0}=0.2$, without additional tuning.
-Importantly, we did not assume $f_{i}$ is $L$-smooth (or even $L_{i}$-smooth), a standard assumption in nonconvex optimization and FL literature Ward et al.
-(2020); Koloskova et al.
-(2020); Li et al.
-(2020); Reddi et al.
-(2021); instead, we can obtain a smaller quantity, $\tilde{L}$, through our analysis in a similar vein to (5); for space limitation, we defer the details and the proof to Appendix A.
+We remind again that the conditions on $\gamma$ and $\eta_{0}$ are only required for our theory.
+[^3] In all experiments, we use the default settings $\gamma=2$, $\eta_{0}=0.2$, and $\theta_{0}=1$ without additional tuning.
+Importantly, we did not assume $f_{i}$ is globally $L$-smooth, a standard assumption in nonconvex optimization and FL literature (Ward et al., 2020; Koloskova et al., 2020; Li et al., 2020; Reddi et al., 2021); instead, we can obtain a smaller quantity, $\tilde{L}$, through our analysis; for space limitation, we defer the details and the proof to Appendix A.
 4 Experimental Setup and Results
+We now introduce the experimental setup and discuss the results. Our implementation can be found in https://github.com/jlylekim/auto-tuned-FL.
 Datasets and models.
-To test the performance of different client optimizers in various settings, we evaluate on four datasets commonly used in FL scenarios: MNIST, FMNIST, CIFAR-10, and CIFAR-100 Krizhevsky et al.
-(2009).
+We consider image classification and text classification tasks.
+We use four datasets for image classification: MNIST, FMNIST, CIFAR-10, and CIFAR-100 (Krizhevsky et al., 2009).
 For MNIST and FMNIST, we train a shallow CNN with two convolutional and two fully-connected layers, followed by dropout and ReLU activations.
-For CIFAR-10, we train a ResNet-18 He et al.
-(2016).
-For CIFAR-100, we train both a ResNet-18 and a ResNet-50 to study the effect of changing the model architecture.
-For each dataset, we create a federated version by randomly partitioning the training data among $100$ clients, with each client getting $500$ examples.
-[^4] All the datasets we consider have $50,000$ training samples, leading to $500$ samples distributed per client.
-To control the level of “non-iidness”, we apply latent Dirichlet allocation (LDA) over the labels following Hsu et al.
-(2019), where each client has an associated multinomial distribution over the labels from which its examples are drawn.
-That is, the local training examples of each client has been drawn from a categorical distribution over $N$ classes, parameterized by $\mathbf{q}$, such that $\mathbf{q}\sim\text{Dir}(\alpha\mathbf{p}),$ where $\mathbf{p}$ is a prior distribution over $N$ classes, with $\alpha>0$ being the concentration parameter.
-We vary $\alpha\in\{0.01,0.1,1\}$, where bigger $\alpha$ indicates settings closer to i.i.d.
-scenarios.
-The class distribution of different $\alpha$’s is visualized in Figure 5 in Appendix B.
+For CIFAR-10, we train a ResNet-18 (He et al., 2016).
+For CIFAR-100, we train both ResNet-18 and ResNet-50 to study the effect of changing the model architecture.
+For text classification, we use two datasets: DBpedia and AGnews datasets (Zhang et al., 2015), and train a DistillBERT (Sanh et al., 2019) for classification.
+For each dataset, we create a federated version by randomly partitioning the training data among $100$ clients ($50$ for text classification), with each client getting $500$ examples.
+To control the level of non-iidness, we apply latent Dirichlet allocation (LDA) over the labels following Hsu et al.
+(2019), where the degree class heterogeneity can be parameterized by the Dirichlet concentration parameter $\alpha.$ The class distribution of different $\alpha$’s is visualized in Figure 8 in Appendix B.9.
 FL setup and optimizers.
-For all cases, we fix the number of clients to be $100,$ and randomly sample $10\%$ among the $100$ clients.
-Similarly to Reddi et al.
-(2021); Li et al.
-(2020), we perform $E$ local epochs of training over each client’s dataset, and we utilize mini-batch gradients of size $b=64$, leading to $K\approx\lfloor\tfrac{E\cdot 500}{64}\rfloor$ local gradient steps; we use $E=1$ for all settings.
-For client optimizers, we compare stochastic gradient descent (SGD), SGD with momentum (SGDM), adaptive methods including Adam Kingma and Ba (2014), Adagrad Duchi et al.
-(2011), SGD with stochastic Polyak step size (SPS) Loizou et al.
-(2021), and our proposed method: $\Delta$-SGD in Algorithm 1.
-As our focus is on client adaptivity, we only present the results using simple FedAvg McMahan et al.
-(2017) as the server optimizer.
-[…]
-Implementation.
-We use Pytorch Paszke et al.
-(2019) to implement all experiments and run on eight P100 GPUs.
-Our implementation will be available online after the anonymous review process for reproducibility.
-Figure 2: The effect of stronger heterogeneity on different client optimizers, induced by the Dirichlet concentration parameter $\alpha$.
+We fix the number of clients to be $100$ for the image classification, and $50$ for the text classification; we randomly sample $10\%$ as participating clients.
+We perform $E$ local epochs of training over each client’s dataset.
+We utilize mini-batch gradients of size $b=64$ for the image classification, and $b=16$ for the text classification, leading to $K\approx\lfloor\tfrac{E\cdot 500}{b}\rfloor$ local gradient steps; we use $E=1$ for all settings.
+For client optimizers, we compare stochastic gradient descent (SGD), SGD with momentum (SGDM), adaptive methods including Adam (Kingma & Ba, 2014), Adagrad (Duchi et al., 2011), SGD with stochastic Polyak step size (SPS) (Loizou et al., 2021), and our proposed method: $\Delta$-SGD in Algorithm 1.
+As our focus is on client adaptivity, we mostly present the results using FedAvg (McMahan et al., 2017) as the server optimizer; additional results using FedAdam can be found in Appendix B.4.
+Hyperparameters.
+For each optimizer, we perform a grid search of learning rates on a single task: CIFAR-10 classification trained with a ResNet-18, with Dirichlet concentration parameter $\alpha=0.1$; for the rest of the settings, we use the same learning rates.
+For SGD, we perform a grid search with $\eta\in\{0.01,0.05,0.1,0.5\}$.
+For SGDM, we use the same grid for $\eta$ and use momentum parameter $\beta=0.9$.
+To properly account for the SGD(M) fine-tuning typically done in practice, we also test dividing the step size by 10 after $50\%$, and again by 10 after $75\%$ of the total training rounds (LR decay).
+For Adam and Adagrad, we grid search with $\eta\in\{0.001,0.01,0.1\}.$ For SPS, we use the default setting of the official implementation.
+[^4] I.e., we use $f_{i}^{\star}=0$, and $c=0.5,$ for the SPS step size: $\frac{f_{i}(x)-f_{i}^{\star}}{c\|\nabla f_{i}(x)\|^{2}}$.
+The official implementation can be found in https://github.com/IssamLaradji/sps.
+For $\Delta$-SGD, we append $\delta$ in front of the second condition: $\sqrt{1+\delta\theta_{t,k-1}^{i}}\eta_{t,k-1}^{i}$ following Malitsky & Mishchenko (2020), and use $\delta=0.1$ for all experiments.
+[^5] We also demonstrate in Appendix B.1 that $\delta$ has very little impact on the final accuracy of $\Delta$-SGD.
+Finally, for the number of rounds $T$, we use $500$ for MNIST, $1000$ for FMNIST, $2000$ for CIFAR-10 and CIFAR-100, and $100$ for the text classification tasks.
 4.1 Results
-Results of the experiments detailed in the previous subsection are presented in Table 1.
-Below, we make comprehensive remarks, depending on: $i)$ the level of heterogeneity induced by different $\alpha$’s of the Dirichlet concentration parameter, $ii)$ the effect of changing the model architecture, and $iii)$ the effect of changing the dataset, and how these affect the performance of different client optimizers.
+We clarify again that the best step size for each client optimizer was tuned via grid-search for a single task: CIFAR-10 classification trained with a ResNet-18, with Dirichlet concentration parameter $\alpha=0.1$.
+We then intentionally use the same step size in all other tasks to highlight two points: $i)$ $\Delta$-SGD works well without any tuning across different datasets, model architectures, and degrees of heterogeneity; $ii)$ other optimizers perform suboptimally without additional tuning.
 Non-iidness Optimizer Dataset / Model
@@
 CNN CNN ResNet-18 ResNet-18 ResNet-50
-$\alpha=1$ SGD 98.3${}_{\downarrow(0.2)}$ 86.5${}_{\downarrow(0.8)}$ 87.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 57.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.2)}}$ 53.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.8)}}$
-SGD ($\downarrow$) 97.8${}_{\downarrow(0.7)}$ 86.3${}_{\downarrow(1.0)}$ 87.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.0)}}$ 61.9${}_{\downarrow(0.0)}$ 60.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.9)}}$
-SGDM 98.5${}_{\downarrow(0.0)}$ 85.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 88.7${}_{\downarrow(1.1)}$ 58.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.1)}}$ 60.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.3)}}$
-SGDM ($\downarrow$) 98.4${}_{\downarrow(0.1)}$ 87.2${}_{\downarrow(0.1)}$ 89.3${}_{\downarrow(0.5)}$ 61.4${}_{\downarrow(0.5)}$ 63.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.5)}}$
-Adam 94.7${}_{{\color[rgb]{0.96,0.29,0.54}{\color[rgb]{0.96,0.29,0.54}\downarrow(3.8)}}}$ 71.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.5)}}$ 89.4${}_{\downarrow(0.4)}$ 55.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.3)}}$ 61.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.4)}}$
-Adagrad 64.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(34.2)}}$ 45.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(41.8)}}$ 86.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.2)}}$ 53.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.4)}}$ 51.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.9)}}$
-SPS 10.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(88.4)}}$ 85.9${}_{\downarrow(1.4)}$ 82.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$ 1.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(60.9)}}$ 50.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.8)}}$
-$\Delta$-SGD 98.4${}_{\downarrow(0.1)}$ 87.3${}_{\downarrow(0.0)}$ 89.8${}_{\downarrow(0.0)}$ 61.5${}_{\downarrow(0.4)}$ 65.8${}_{\downarrow(0.0)}$
-$\alpha=0.1$ SGD 98.1${}_{\downarrow(0.0)}$ 83.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.8)}}$ 72.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.9)}}$ 54.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.7)}}$ 44.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.9)}}$
-SGD ($\downarrow$) 98.0${}_{\downarrow(0.1)}$ 84.7${}_{\downarrow(1.7)}$ 78.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.6)}}$ 59.3${}_{\downarrow(1.8)}$ 48.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.4)}}$
-SGDM 97.6${}_{\downarrow(0.5)}$ 83.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.8)}}$ 79.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.4)}}$ 58.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.3)}}$ 52.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.8)}}$
-SGDM ($\downarrow$) 98.0${}_{\downarrow(0.1)}$ 86.1${}_{\downarrow(0.3)}$ 77.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$ 60.4${}_{\downarrow(0.7)}$ 52.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.3)}}$
-Adam 96.4${}_{\downarrow(1.7)}$ 80.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.0)}}$ 85.0${}_{\downarrow(0.0)}$ 55.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.7)}}$ 58.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.9)}}$
-Adagrad 89.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.2)}}$ 46.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(40.1)}}$ 84.1${}_{\downarrow(0.9)}$ 49.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.5)}}$ 48.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.1)}}$
-SPS 96.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 85.0${}_{\downarrow(1.4)}$ 70.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.7)}}$ 42.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(18.9)}}$ 42.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(21.9)}}$
-$\Delta$-SGD 98.1${}_{\downarrow(0.0)}$ 86.4${}_{\downarrow(0.0)}$ 84.5${}_{\downarrow(0.5)}$ 61.1${}_{\downarrow(0.0)}$ 64.1${}_{\downarrow(0.0)}$
-$\alpha=0.01$ SGD 96.8${}_{\downarrow(0.7)}$ 79.0${}_{\downarrow(1.2)}$ 22.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.3)}}$ 30.5${}_{\downarrow(1.3)}$ 24.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$
-SGD ($\downarrow$) 97.2${}_{\downarrow(0.3)}$ 79.3${}_{\downarrow(0.9)}$ 33.9${}_{\downarrow(0.0)}$ 30.3${}_{\downarrow(1.5)}$ 24.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.8)}}$
-SGDM 77.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.6)}}$ 75.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.5)}}$ 28.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.5)}}$ 24.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.0)}}$ 22.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.4)}}$
-SGDM ($\downarrow$) 94.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.5)}}$ 79.5${}_{\downarrow(0.7)}$ 29.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.9)}}$ 20.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(10.9)}}$ 14.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.7)}}$
-Adam 80.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.7)}}$ 60.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.6)}}$ 22.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.8)}}$ 18.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.6)}}$ 22.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.8)}}$
-Adagrad 72.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(25.1)}}$ 45.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(34.3)}}$ 12.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(21.4)}}$ 25.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.0)}}$ 22.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.2)}}$
-SPS 69.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(27.8)}}$ 44.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(36.2)}}$ 21.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.4)}}$ 22.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.8)}}$ 17.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.0)}}$
-$\Delta$-SGD 97.5${}_{\downarrow(0.0)}$ 80.2${}_{\downarrow(0.0)}$ 31.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.3)}}$ 31.8${}_{\downarrow(0.0)}$ 31.4${}_{\downarrow(0.0)}$
+$\alpha=1$ SGD 98.3↓(0.2) 86.5↓(0.8) 87.7↓(2.1) 57.7↓(4.2) 53.0↓(12.8)
+SGD ($\downarrow$) 97.8↓(0.7) 86.3↓(1.0) 87.8↓(2.0) 61.9↓(0.0) 60.9↓(4.9)
+SGDM 98.5↓(0.0) 85.2↓(2.1) 88.7↓(1.1) 58.8↓(3.1) 60.5↓(5.3)
+SGDM ($\downarrow$) 98.4↓(0.1) 87.2↓(0.1) 89.3↓(0.5) 61.4↓(0.5) 63.3↓(2.5)
+Adam 94.7↓(3.8) 71.8↓(15.5) 89.4↓(0.4) 55.6↓(6.3) 61.4↓(4.4)
+Adagrad 64.3↓(34.2) 45.5↓(41.8) 86.6↓(3.2) 53.5↓(8.4) 51.9↓(13.9)
+SPS 10.1↓(88.4) 85.9↓(1.4) 82.7↓(7.1) 1.0↓(60.9) 50.0↓(15.8)
+$\Delta$-SGD 98.4↓(0.1) 87.3↓(0.0) 89.8↓(0.0) 61.5↓(0.4) 65.8↓(0.0)
+$\alpha=0.1$ SGD 98.1↓(0.0) 83.6↓(2.8) 72.1↓(12.9) 54.4↓(6.7) 44.2↓(19.9)
+SGD ($\downarrow$) 98.0↓(0.1) 84.7↓(1.7) 78.4↓(6.6) 59.3↓(1.8) 48.7↓(15.4)
+SGDM 97.6↓(0.5) 83.6↓(2.8) 79.6↓(5.4) 58.8↓(2.3) 52.3↓(11.8)
+SGDM ($\downarrow$) 98.0↓(0.1) 86.1↓(0.3) 77.9↓(7.1) 60.4↓(0.7) 52.8↓(11.3)
+Adam 96.4↓(1.7) 80.4↓(6.0) 85.0↓(0.0) 55.4↓(5.7) 58.2↓(5.9)
+Adagrad 89.9↓(8.2) 46.3↓(40.1) 84.1↓(0.9) 49.6↓(11.5) 48.0↓(16.1)
+SPS 96.0↓(2.1) 85.0↓(1.4) 70.3↓(14.7) 42.2↓(18.9) 42.2↓(21.9)
+$\Delta$-SGD 98.1↓(0.0) 86.4↓(0.0) 84.5↓(0.5) 61.1↓(0.0) 64.1↓(0.0)
+$\alpha=0.01$ SGD 96.8↓(0.7) 79.0↓(1.2) 22.6↓(11.3) 30.5↓(1.3) 24.3↓(7.1)
+SGD ($\downarrow$) 97.2↓(0.3) 79.3↓(0.9) 33.9↓(0.0) 30.3↓(1.5) 24.6↓(6.8)
+SGDM 77.9↓(19.6) 75.7↓(4.5) 28.4↓(5.5) 24.8↓(7.0) 22.0↓(9.4)
+SGDM ($\downarrow$) 94.0↓(3.5) 79.5↓(0.7) 29.0↓(4.9) 20.9↓(10.9) 14.7↓(16.7)
+Adam 80.8↓(16.7) 60.6↓(19.6) 22.1↓(11.8) 18.2↓(13.6) 22.6↓(8.8)
+Adagrad 72.4↓(25.1) 45.9↓(34.3) 12.5↓(21.4) 25.8↓(6.0) 22.2↓(9.2)
+SPS 69.7↓(27.8) 44.0↓(36.2) 21.5↓(12.4) 22.0↓(9.8) 17.4↓(14.0)
+$\Delta$-SGD 97.5↓(0.0) 80.2↓(0.0) 31.6↓(2.3) 31.8↓(0.0) 31.4↓(0.0)
 Table 1: Experimental results based on the settings detailed in Section 4.
-Performance difference within 0.5% of the best result for each task are shown in bold.
-Subscripts${}_{\downarrow(x.x)}$ is the performance difference from the best result and is highlighted in pink when the difference is bigger than 2%.
-The symbol ($\downarrow$) appended to SGD and SGDM indicates step-wise learning rate decay, where the step sizes are divided by $10$ after 50%, and another by $10$ after 75% of the total rounds.
+Best accuracy ($\pm$ 0.5%) for each task are shown in bold.
+Subscripts↓(x.x) is the performance difference from the best result, and is highlighted in pink when it is bigger than 2%.
+The down-arrow symbol ($\downarrow$) indicates step-wise learning rate decay, where the step sizes are divided by $10$ after 50%, and another by $10$ after 75% of the total rounds.
+Figure 2: The effect of stronger heterogeneity on different client optimizers, induced by the Dirichlet concentration parameter $\alpha\in\{0.01,0.1,1\}$.
+$\Delta$-SGD remains robust performance in all cases, whereas other methods show significant performance degradation when changing the level of heterogeneity $\alpha$, or when changing the setting (model/architecture).
 Changing the level of non-iidness.
@@
 Three illustrative cases are visualized in Figure 2.
-We remind that the hyperparameters are tuned for the case of CIFAR-10 trained with ResNet-18, with $\alpha=0.1$ (Figure 2(A)).
+We remind that the step sizes are tuned for the task in Figure 2(A).
 For this task, with $\alpha=1$ (i.e., closer to iid), all methods perform better than $\alpha=0.1,$ as expected.
-With $\alpha=0.01,$ which is highly non-iid (c.f., Figure 5 in Appendix B), we see a significant drop in performance for all methods.
+With $\alpha=0.01,$ which is highly non-iid (c.f., Figure 8 in Appendix B.9), we see a significant drop in performance for all methods.
 SGD with LR decay and $\Delta$-SGD perform the best, while adaptive methods like Adam ($85\%\rightarrow 22.1\%$) and Adagrad ($84.1\%\rightarrow 12.5\%$) degrade noticeably more than other methods.
-For FMNIST trained with CNN (Figure 2(B)), which can be considered as an easier problem compared to CIFAR-10 classification, the performance degradation with varying $\alpha$’s is much milder.
-Interestingly, adaptive methods like Adam, Adagrad, and SPS perform much worse than other methods, indicating their inflexibility when applied to diverse datasets and model architectures, which often is worse than simple SGD-based methods; we further investigate this phenomenon in more detail in the next remark.
-A similar trend can be seen in MNIST classification, where SGD-based methods work relatively well, except for SGDM with $\alpha=0.01$: without the LR decay, SGDM only achieves around 78% accuracy, while SGD without LR decay still achieves over $96\%$ accuracy (c.f., Figure 1(A)).
-Even with the LR decay, SGDM achieves noticeably worse performance.
-While this may seem like an “easy fix” in hindsight, it indicates that in a highly non-iid setting, one should be extremely careful in tuning SGD-based methods, even for easy datasets like MNIST.
-Interestingly, SPS performs particularly poorly, merely achieving 10% for MNIST with $\alpha=1,$ while achieving 96% for “harder” case of $\alpha=0.1,$ and again dropping to 69.7% with $\alpha=0.01$; a possible explanation is, with simple CNN, the model is not over-parameterized enough to have $f_{i}^{\star}=0,$ the default estimate we use on which the performance of SPS crucially rely Hazan and Kakade (2019).
-Adagrad similarly performs quite poorly.
-Lastly, in Figure 2(C), results for CIFAR-100 classification trained with ResNet-50 are illustrated.
+Figure 2(B) shows the results of training FMNIST with a CNN.
+This problem is generally considered easier than CIFAR-10 classification.
+The experiments show that the performance degradation with varying $\alpha$’s is much milder in this case.
+Interestingly, adaptive methods such as Adam, Adagrad, and SPS perform much worse than other methods
+Lastly, in Figure 2(C), results for CIFAR-100 classification trained with a ResNet-50 are illustrated.
 $\Delta$-SGD exhibits superior performance in all cases of $\alpha$.
@@
 Other methods, including SGD with and without momentum/LR decay, Adagrad, and SPS, perform much worse than $\Delta$-SGD.
-Changing the model architecture.
+Changing the model architecture/dataset.
 For this remark, let us focus on CIFAR-100 trained on ResNet-18 versus on ResNet-50, with $\alpha=0.1$, illustrated in Figure 3(A).
 SGD and SGDM (both with and without LR decay), Adagrad, and SPS perform worse using ResNet-50 than ResNet-18.
-This is a counter-intuitive behavior, as one would expect to get better accuracy by using a more powerful model.
+This is a counter-intuitive behavior, as one would expect to get better accuracy by using a more powerful model, although the step sizes are tuned on ResNet-18.
 $\Delta$-SGD is an exception: without any additional tuning, $\Delta$-SGD can improve its performance.
 Adam also improves similarly, but the achieved accuracy is significantly worse than that of $\Delta$-SGD.
-While it is true that the performance of $\Delta$-SGD also slightly degrades for $\alpha=0.01,$ it still performs better than the other methods.
-Changing the dataset.
 We now focus on cases where the dataset changes, but the model architecture remains the same.
 We mainly consider two cases, illustrated in Figure 3(B) and (C): When CNN is trained for classifying MNIST versus FMNIST, and when ResNet-18 is trained for CIFAR-10 versus CIFAR-100.
-Aggravating the complexity of fine-tuning SGD(M) for MNIST, one can observe a similar trend in CIFAR-10 trained with ResNet-18.
+From (B), one can infer the difficulty in tuning SGDM: without the LR decay, SGDM only achieves around 78% accuracy for MNIST classification, while SGD without LR decay still achieves over $96\%$ accuracy.
+For FMNIST on the other hand, SGDM performs relatively well, but adaptive methods like Adam, Adagrad, and SPS degrade significantly.
+Aggravating the complexity of fine-tuning SGD(M) for MNIST, one can observe from (C) a similar trend in CIFAR-10 trained with ResNet-18.
 In that case, $\Delta$-SGD does not achieve the best test accuracy (although it is the second best with a pretty big margin with the rest), while SGD with LR decay does.
 However, without LR decay, the accuracy achieved by SGD drops more than 11%.
-Interestingly, for the same setting, when we change the dataset from CIFAR-10 to CIFAR-100, SGD without LR decay (which is not one of the best methods for the CIFAR-10 case) achieves the second-best performance (after $\Delta$-SGD).
-On the other hand, SGDM with LR decay, which achieves the 3rd best performance for CIFAR-10 achieves very poor performance (more than 10% less than the best) for CIFAR-100.
-Finally, SGD with decay, which performs the best in the CIFAR-10 case, turns out to be the 3rd best algorithm when we change the dataset to CIFAR-100.
-Again, adaptive methods like Adam, Adagrad, and SPS perform quite poorly in this case.
+Again, adaptive methods like Adam, Adagrad, and SPS perform suboptimally in this case.
 Figure 3: The effect of changing the dataset and the model architecture on different client optimizers.
+$\Delta$-SGD remains superior performance without additional tuning when model or dataset changes, whereas other methods often degrades in performance.
 (A): CIFAR-100 trained with ResNet-18 versus Resnet-50 ($\alpha=0.1$), (B): MNIST versus FMNIST trained with CNN ($\alpha=0.01$), (C): CIFAR-10 versus CIFAR-100 trained with ResNet-18 ($\alpha=0.01$).
+Changing the domain: text classification.
+In Table 4(a), we make a bigger leap, and test the performance of different client optimizers for text classification tasks.
+We use the DBpedia and AGnews datasets (Zhang et al., 2015), and train a DistillBERT (Sanh et al., 2019) for classification.
+Again, the best step size for each client optimizer was tuned for the CIFAR-10 image classification trained with a ResNet-18, for $\alpha=0.1$.
+Thus, the type of the dataset changed completely from image to text, as well as the model architecture: from vision to language.
+Not surprisingly, four methods (SGDM, SGDM($\downarrow$), Adam, and Adagrad) ended up learning nothing (i.e., the achieved accuracy is 1/(# of labels)), indicating the fine-tuned step sizes for these methods using CIFAR-10 classification task was too large for the considered text classification tasks.
+$\Delta$-SGD, on the other hand, still achieves competitive accuracies without additional tuning, thanks to the locality adaptive step size.
+Interestingly, SGD, SGD($\downarrow$), and SPS worked well for the text classification tasks, which is contrary to their suboptimal performances for image classification tasks in Table 1.
+Changing the loss function.
+In Table 4(b), we illustrate how $\Delta$-SGD performs when combined with a different loss function, such as FedProx (Li et al., 2020).
+FedProx aims to address heterogeneity; we thus used the lowest concentration parameter, $\alpha=0.01$.
+We only present a subset of results: CIFAR-10(100) classification with ResNet-18(50) (additional results in Appendix B.6).
+The results show that $\Delta$-SGD again outperforms all other methods with a significant margin: $\Delta$-SGD is not only the best, but also the performance difference is at least 2.5%, and as large as 26.9%.
+Yet, compared to the original setting without the proximal term in Table 1, it is unclear whether or not the additional proximal term helps.
+For $\Delta$-SGD, the accuracy increased for CIFAR-10 with ResNet-18, but slightly got worse for CIFAR-100 with ResNet-50; a similar story applies to other methods.
+Text classification Dataset / Model
+$\alpha=1$ Agnews Dbpedia
+Optimizer DistillBERT
+SGD 91.1↓(0.5) 96.0↓(2.9)
+SGD ($\downarrow$)↓ ) 91.6↓(0.0) 98.7↓(0.2)
+SGDM 25.0↓(66.6) 7.1↓(91.8)
+SGDM ($\downarrow$)↓ ) 25.0↓(66.6) 7.1↓(91.8)
+Adam 25.0↓(66.6) 7.1↓(91.8)
+Adagrad 25.0↓(66.6) 7.1↓(91.8)
+SPS 91.5↓(0.1) 98.9↓(0.0)
+$\Delta$-SGD 90.7↓(0.9) 98.6↓(0.3)
+FedProx Dataset / Model
+$\alpha=0.01$ CIFAR-10 CIFAR-100
+Optimizer ResNet-18 ResNet-50
+SGD 20.0↓(13.8) 25.2↓(5.9)
+SGD ($\downarrow$)↓ ) 31.3↓(2.5) 20.2↓(10.8)
+SGDM 29.3↓(4.4) 23.8↓(7.2)
+SGDM ($\downarrow$)↓ ) 25.3↓(8.5) 15.0↓(16.0)
+Adam 28.1↓(5.7) 22.6↓(8.4)
+Adagrad 19.3↓(14.5) 4.1↓(26.9)
+SPS 27.6↓(6.2) 16.5↓(14.5)
+$\Delta$-SGD 33.8↓(0.0) 31.0↓(0.0)
+Table 2: Additional experiments for a different domain (a), and a different loss function (b).
+Other findings.
+We present all the omitted theorems and proofs in Appendix A, and additional empirical findings in Appendix B.
+Specifically, we demonstrate: the robustness of the performance of $\Delta$-SGD (Appendix B.1), the effect of the different numbers of local iterations (Appendix B.2), additional experiments when the size of local dataset is different per client (Appendix B.3), additional experimental results using FedAdam (Reddi et al., 2021) (Appendix B.4), additional experiential results using MOON (Appendix B.5) and FedProx (Appendix B.6) loss functions, and three independent trials for a subset of tasks (Appendix B.7).
+We also visualize the step size conditions for $\Delta$-SGD (Appendix B.8) as well as the degree of heterogeneity (Appendix B.9).
 5 Conclusion
-In this work, we propose $\Delta$-SGD, a distributed SGD scheme equipped with an adaptive step size that enables each client to use its step size and adapts to the local smoothness of the function each client is optimizing.
-We prove the convergence of $\Delta$-SGD in the general nonconvex setting and present extensive empirical results, where the superiority of $\Delta$-SGD is shown in various scenarios without any tuning.
-For future works, extending $\Delta$-SGD to a coordinate-wise step size in spirit of Duchi et al.
-(2011); Kingma and Ba (2014), as well as enabling asynchronous updates Assran et al.
-(2020); Toghani et al.
-(2022); Nguyen et al.
-(2022) could be interesting directions.
-Broader Impacts and Limitations.
-We believe our proposed method, $\Delta$-SGD, can potentially impact reducing the amount of tuning required in training FL models.
-By now, it is widely believed that FL systems do not “dodge the bullet” of high carbon emission in AI/ML: as noted in a recent study Yousefpour et al.
-(2023), “compute on client devices, and the communication between the clients and the server are responsible for the majority of FL’s overall carbon emissions (97%).” The lack of principled hyperparameter tuning only exacerbates this problem.
-Based on the results presented in Section 4, $\Delta$-SGD achieves superior performance while robust to different settings.
-In terms of limitations, our assumption on Theorem 1, especially Assumption 1c, as well as the conditions on $\gamma$ and $\eta_{0}$ are quite strong, and does not truly capture the strength of $\Delta$-SGD.
-Moreover, due to non-optimal implementation, $\Delta$-SGD takes about 1.3×1.3$\times$1.3 × more wall-clock time compared to SGD, which can be improved with more efficient implementation.
+In this work, we proposed $\Delta$-SGD, a distributed SGD scheme equipped with an adaptive step size that enables each client to use its own step size and adapts to the local smoothness of the function each client is optimizing.
+We proved the convergence of $\Delta$-SGD in the general nonconvex setting and presented extensive empirical results, where the superiority of $\Delta$-SGD is shown in various scenarios without any tuning.
+For future works, extending $\Delta$-SGD to a coordinate-wise step size in the spirit of Duchi et al.
+(2011); Kingma & Ba (2014), applying the step size to more complex algorithms like (local) factored gradient descent (Kim et al., 2022a), and enabling asynchronous updates (Assran et al., 2020; Toghani et al., 2022; Nguyen et al., 2022) could be interesting directions.
+Acknowledgments
+This work is supported by NSF FET: Small No.
+1907936, NSF MLWiNS CNS No.
+2003137 (in collaboration with Intel), NSF CMMI No.
+2037545, NSF CAREER award No.
+2145629, NSF CIF No.
+2008555, Rice InterDisciplinary Excellence Award (IDEA), NSF CCF No.
+2211815, NSF DEB No.
+2213568, and Lodieska Stockbridge Vaughn Fellowship.
 Supplementary Materials for
@@
     • In Section A.1, we provide the proof of Algorithm 1, under Assumption 1.
-    • In Section A.2, we provide the proof of Algorithm, 1, with some additional assumptions, including that $f_{i}$ is convex for all $i\in[m]$.
-    • In Section A.3, we provide a simple proof in the centralized nonconvex setting, which was an open question posed in Malitsky and Mishchenko (2020).
+    • In Section A.2, we provide the proof of Algorithm, 1, with some modified assumptions, including that $f_{i}$ is convex for all $i\in[m]$.
   • In Appendix B, we provide additional experiments, as well as miscellaneous plots that were missing in the main text due to the space constraints. Specifically:
     • In Section B.1, we provide some perspectives on the ease of tuning of $\Delta$-SGD in Algorithm 1.
-    • In Section B.2, we provide illustration of the different level of heterogeneity induced by the Dirichlet concentration parameter $\alpha$.
     • In Section B.2, we provide some perspectives on how $\Delta$-SGD performs using different number of local iterations.
     • In Section B.3, we provide additional experimental results where each client has different number of samples as local dataset.
+    • In Section B.4, we provide additional experimental results using FedAdam (Reddi et al., 2021) server-side adaptive method.
+    • In Section B.5, we provide additional experimental results using MOON (Li et al., 2021) loss function.
+    • In Section B.6, we provide additional experiemtns using FedProx (Li et al., 2020) loss function.
+    • In Section B.7, we perform three independent trials for a subset of tasks for all client-optimizers.
+We plot the average and the standard deviations in Figure 6, and report those values in Table 7.
+    • In Section B.8, we visualize the step size conditions for $\Delta$-SGD.
+    • In Section B.9, we provide illustration of the different level of heterogeneity induced by the Dirichlet concentration parameter $\alpha$.
 Appendix A Missing Proofs
-We first introduce the following two inequalities which will be used in the following proofs.
+We first introduce the following two inequalities, which will be used in the subsequent proofs:
 $\langle x_{i},x_{j}\rangle$ $\leqslant\frac{1}{2}\|x_{i}\|^{2}+\frac{1}{2}\|x_{j}\|^{2},\quad\text{and}$ (7)
 $\Big{\|}\sum_{i=1}^{m}x_{i}\Big{\|}^{2}$ $\leqslant m\sum_{i=1}^{m}\Big{\|}x_{i}\big{\|}^{2},$ (8)
-for any set of $m$ vectors $\{x_{i}\}_{i{=}1}^{m}$ such that $x_{i}\in\mathbb{R}^{d}$.
-We also provide the following lemma, which establishes $\tilde{L}$-smoothness based on local smoothness, without assuming the global $L$-smoothness.
+for any set of $m$ vectors $\{x_{i}\}_{i{=}1}^{m}$ with $x_{i}\in\mathbb{R}^{d}$.
+We also provide the following lemma, which establishes $\tilde{L}$-smoothness based on the local smoothness.
 Lemma 2.
-For locally smooth function $f_{i}$, the following holds for the points that Algorithm 1 visits:
-$f_{i}(x_{t})\leqslant f_{i}(x_{t-1})+\langle\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$ (9)
+For locally smooth function $f_{i}$, the following inequality holds:
+$f_{i}(y)\leqslant f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{\tilde{L}_{i,t}}{2}\|y-x\|^{2}\quad\forall x,y\in\mathcal{S}_{i,t},$ (9)
+where $\mathcal{S}_{i,t}:=\mathrm{conv}(\{x_{t,k}^{i}\}_{k=1}^{K}\cup\{x_{t},x_{t+1}\})$.
 Proof.
-By the definition of $\eta_{t}^{i},$ we have
-$\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1})\|\leqslant\frac{1}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|$
+First, notice that by Assumptions (1a) and (1b), stochastic gradients are also bounded for all $x\in\mathbb{R}^{d}.$ Therefore, the local iterates, i.e., $\{x_{t,k}^{i}\}_{k=0}^{K}$ remain bounded.
+Hence, the set $\mathcal{S}_{i,t}:=\mathrm{conv}(\{x_{t,k}^{i}\}_{k=1}^{K}\cup\{x_{t},x_{t+1}\})$, where $\mathrm{conv}(\cdot)$ denotes the convex hull, is bounded.
+Due to the fact that $f_{i}$ is locally smooth, $\nabla f$ is Lipschitz continuous on bounded sets.
+This implies there exists a positive constant $\tilde{L}_{i,t}$ such that,
+$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leqslant\tilde{L}_{i,t}\|y-x\|\quad\forall x,y\in\mathcal{S}_{i,t}.$ (10)
 Thus, we have
-$\langle\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle$ $\leqslant\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1})\|\cdot\|x_{t}-x_{t-1}\|$
-$\leqslant\frac{1}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
-Let $g(k)=f_{i}(x_{t-1}+k(x_{t}-x_{t-1})).$ Then, we have
-$g^{\prime}(k)$ $=\langle\nabla f_{i}(x_{t-1}+k(x_{t}-x_{t-1})),x_{t}-x_{t-1}\rangle$
-$g^{\prime}(k)-g^{\prime}(0)$ $=\langle\nabla f_{i}(x_{t-1}+k(x_{t}-x_{t-1})-\nabla f_{i}(x_{t}),\frac{k}{k}(x_{t}-x{t-1})\rangle$
-$\leqslant\frac{k}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
+$\langle\nabla f_{i}(y)-\nabla f_{i}(x),y-x\rangle$ $\leqslant\|\nabla f_{i}(y)-\nabla f_{i}(x)\|\cdot\|y-x\|$
+$\leqslant\tilde{L}_{i,t}\|y-x\|^{2},\quad\forall x,y\in\mathcal{S}_{i,t}.$
+Let $g(\tau)=f_{i}(x+\tau(y-x)).$ Then, we have
+$g^{\prime}(\tau)$ $=\langle\nabla f_{i}(x+\tau(y-x)),y-x\rangle$
+$g^{\prime}(\tau)-g^{\prime}(0)$ $=\langle\nabla f_{i}(x+\tau(y-x))-\nabla f_{i}(x),y-x\rangle$
+$\leqslant\|\nabla f_{i}(x+\tau(y-x))-\nabla f_{i}(x)\|\cdot\|y-x\|$
+$\leqslant\tau\tilde{L}_{i,t}\|y-x\|^{2}.$
 Using the mean value theorem, we have
@@
 […]
-$\leqslant g(0)+g^{\prime}(0)+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}$
-$\leqslant f_{i}(x_{t-1})+\langle\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
-We also establish the averaged version of local $\tilde{L}$-smoothness. First, we re-write the statement of Lemma 2 as follows:
-$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}\leqslant\tilde{L}_{i}\|x-y\|^{2},$ (10)
-where $\tilde{L}_{i}$ is $\max_{t}\Big{(}\frac{1}{2\eta_{t}^{i}}\Big{)}$ for all $x,y\in\text{conv}\{x_{0},x_{1},\dots\}$, and $\text{conv}\{\cdots\}$ denotes the convex hull of the arguments Boyd et al.
-(2004).
-Now, given (10),
+$\leqslant g(0)+g^{\prime}(0)+\frac{\tilde{L}_{i,t}}{2}\|y-x\|^{2}$
+$\leqslant f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{\tilde{L}_{i,t}}{2}\|y-x\|^{2}.$
+Now, we define $\tilde{L}_{i}:=\max_{t}\tilde{L}_{i,t}$.
+Given this definition, we relate the local-smoothness constant $\tilde{L}$ of the average function $f$ with (individual) local-smoothness constants $\tilde{L}_{i}$ in the below lemma.
 Lemma 3.
-The average of $\tilde{L}$-local smooth functions $f_{i}$ are at least $\tilde{L}$-local smooth, where $\tilde{L}=\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i}$.
+The average of $\tilde{L}_{i}$-local smooth functions $f_{i}$ are at least $(\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i})$-local smooth.
 Proof.
@@
 $\leqslant\frac{1}{m}\sum_{i=1}^{m}\big{\|}\nabla f_{i}(x)-\nabla f_{i}(y)\big{\|}$
-$\leqslant\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i}\|x-y\|$
-$=\tilde{L}\|x-y\|.$ (11)
+$\leqslant\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i}\|x-y\|$ (11)
+$\leqslant\tilde{L}\|x-y\|,\quad\text{where}\quad\tilde{L}=\max_{i,t}\tilde{L}_{i,t}.$ (12)
+Technically, (11) can provide a slightly tighter bound for Theorem 1, but (12) simplifies the final expression. We also have the following form:
+$f(x_{t+1})\leqslant f(x_{t})+\langle\nabla f(x_{t}),x_{t+1}-x_{t}\rangle+\frac{\tilde{L}}{2}\|x_{t+1}-x_{t}\|^{2},$
+which can be obtained from Lemma 2. ∎
 A.1 Proof of Algorithm 1 under Assumption 1
@@
 where we denoted with $x_{t+1}$ as the server parameter at round $t+1$, which is the average of the local parameters $x_{t,k}^{i}$ for all client $\forall~{}i\in[m]$ after running $k$ local iterations for $\forall~{}k\in[K]$.
-We start from the smoothness of $f(\cdot)$, which we emphasize again that is obtained via Lemma 3, i.e., as a byproduct of our analysis, and we do not assume that $f_{i}$’s are globally $L$-smooth.
-Also note that Lemma 3 holds for $\text{conv}\{x_{0},x_{1},\dots\}$, i.e., the average of the local parameters visited by Algorithm 1.
+We also recall that we use ~$\tilde{\nabla}f_{i}(x)=\frac{1}{|\mathcal{B}|}\sum_{z\in\mathcal{B}}\nabla F_{i}(x,z)$ to denote the stochastic gradients with batch size $|\mathcal{B}|=b$.
+Note that whenever we use this shorthand notation in the theory, we imply that we draw a new independent batch of samples to compute the stochastic gradients.
+We start from the smoothness of $f(\cdot)$, which we note again is obtained via Lemma 3, i.e., as a byproduct of our analysis, and we do not assume that $f_{i}$’s are globally $L$-smooth.
+Also note that Lemma 3 holds for $\{x_{0},x_{1},\dots\}$, i.e., the average of the local parameters visited by Algorithm 1.
 Proceeding, we have
 $f(x_{t+1})\leqslant f(x_{t})-\Big{\langle}\nabla f(x_{t}),\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\rangle}+\frac{\tilde{L}}{2}\Big{\|}\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\|}^{2}.$
-Taking expectation from the expression above, we get
+Taking expectation from the expression above, where we overload the notation such that the expectation is both with respect to the client sampling randomness (indexed with $i$) as well as the data sampling randomness (indexed with $z$), that is $\mathbb{E}(\cdot):=\mathbb{E}_{\mathcal{F}_{t}}\left[\mathbb{E}_{i}\left[\mathbb{E}_{z}[\cdot\mid\mathcal{F}_{t},i]|\mathcal{F}_{t}\right]\right],$ where $\mathcal{F}_{t}$ is the natural filtration, we have:
 $\mathbb{E}f(x_{t+1})$
@@
 $\hskip 113.81102pt+\frac{\tilde{L}}{2m}\sum_{i=1}^{m}\,\mathbb{E}\Big{\|}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\tilde{\nabla}f_{i}(x_{t,k}^{i})\Big{\|}^{2},$ (15)
-where in the equality, we added and subtracted $\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\nabla f(x_{t})$ in the inner product term, and also replaced $D_{t}=\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}$.
+where in the equality, we added and subtracted $\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\nabla f(x_{t})$ in the inner product term, and also replaced $D_{t}:=\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}$.
 Then, using (7) and (13)-(15), we have
@@
 $\qquad\qquad\qquad\qquad\qquad+\nabla f_{i}(x_{t})-\nabla f(x_{t})+\nabla f(x_{t})\Big{)}\Big{\|}^{2}$ (18)
-We now expand the terms in (17), and will bound each seprately, as follows.
+We now expand the terms in (18) and bound each term separately, as follows.
 $\mathbb{E}f(x_{t+1})$ $\leqslant f(x_{t})-\frac{D_{t}}{2}\,\|\nabla f(x_{t})\|^{2}$
@@
 $\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\Big{(}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{)}\Big{\|}^{2}+\,\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
-$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
-$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
+$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,l}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
+$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,l}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
 $\leqslant\frac{2K^{2}\tilde{L}^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}}{D_{t}m}\sum_{i=1}^{m}\Big{(}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}^{2}.$ (21)
@@
 $\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\Big{(}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{)}\Big{\|}^{2}+\,\mathbb{E}\Big{\|}\sum_{\ell=0}^{k-1}\eta_{t,l}^{i}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
-$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
-$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
+$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,l}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
+$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,l}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
 $\leqslant\frac{4K^{2}\tilde{L}^{3}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}}{m}\sum_{i=1}^{m}\Big{(}\sum_{k=0}^{K-1}(\eta_{t,k}^{i})^{2}\Big{)}^{2}.$ (25)
@@
 Finally, by choosing $\gamma=\mathcal{O}\Big{(}\frac{1}{K\sqrt{T}}\Big{)}$, and defining $\Psi_{1}=\max\left\{\frac{\sigma^{2}}{b},f(x_{0})-f(x^{\star})\right\}$ and $\Psi_{2}=\left(\frac{\sigma^{2}}{b}+G^{2}\right)$ with $b=|\mathcal{B}|$ being the batch size, we arrive at the statement in Theorem 1, i.e.,
-$\frac{1}{T}\sum_{t=0}^{T-1}\,\mathbb{E}\Big{\|}\nabla f\Big{(}x_{t}\Big{)}\Big{\|}^{2}$ $\leqslant\mathcal{O}\Big{(}\frac{\Psi_{1}}{\sqrt{T}}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{2}\Psi_{2}}{T}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{3}\Psi_{2}}{\sqrt{T^{3}}}\Big{)}.$
+$\frac{1}{T}\sum_{t=0}^{T-1}\,\mathbb{E}\big{\|}\nabla f(x_{t})\big{\|}^{2}$ $\leqslant\mathcal{O}\Big{(}\frac{\Psi_{1}}{\sqrt{T}}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{2}\Psi_{2}}{T}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{3}\Psi_{2}}{\sqrt{T^{3}}}\Big{)}.$
 A.2 Proof of Algorithm 1 for convex case
-Here, we provide an extension of the proof of (Malitsky and Mishchenko, 2020, Theorem 1) for distributed version.
-Note that the proof we provide here is not exactly the same version of $\Delta$-SGD presented in Algorithm 1; in particular, we assume no local iterations, i.e., $k=1$ for all $i\in[m]$, and every client participate, i.e, $p=1.$
-We start with the extension of (Malitsky and Mishchenko, 2020, Lemma 1), which constructs a Lyapunov function for the distributed objective in (1).
+Here, we provide an extension of the proof of Malitsky & Mishchenko (2020, Theorem 1) for the distributed version.
+Note that the proof we provide here is not exactly the same version of $\Delta$-SGD presented in Algorithm 1; in particular, we assume no local iterations, i.e., $k=1$ for all $i\in[m]$, and every client participate, i.e, $p=1$, without stochastic gradients.
+We start by constructing a Lyapunov function for the distributed objective in (1).
 Lemma 4.
-Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex and twice differentiable. Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$, satisfy the following:
+Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex with locally Lipschitz gradients.
+Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$ with full batch, satisfy the following:
 $\|x_{t+1}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\Big{[}\eta_{t}^{i}(1+\theta_{t}^{i})\big{(}f_{i}(x_{t}^{i})-f_{i}(x^{\star})\big{)}\Big{]}$
@@
 Theorem 5.
-Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex and twice differentiable. Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$, satisfy the following:
-$\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}\leqslant\frac{2DL^{2}}{2Lt+1},$ (38)
+Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex with locally Lipschitz gradients.
+Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$ with full batch, satisfy the following:
+$\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}\leqslant\frac{2D\hat{L}^{2}}{2\hat{L}t+1},$ (39)
 where
@@
 $\alpha_{k}^{i}$ $=\eta_{k}^{i}(1+\theta_{k}^{i})-\eta_{k+1}^{i}\theta_{k+1}^{i}$
-$S_{t}^{i}$ $=\eta_{t}^{i}(1+\theta_{t}^{i})+\sum_{k=1}^{t-1}\alpha_{k}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}.$
-Further, if $x^{\star}$ is any minimum of $f_{i}(x)$ for all $i\in[m]$, (38) implies convergence for problem (1).
+$S_{t}^{i}$ $=\eta_{t}^{i}(1+\theta_{t}^{i})+\sum_{k=1}^{t-1}\alpha_{k}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i},$
+and $\hat{L}$ is a constant. Further, if $x^{\star}$ is any minimum of $f_{i}(x)$ for all $i\in[m]$, (39) implies convergence for problem (1).
 Proof.
@@
 $\leqslant\|x_{1}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{1}^{i}-x_{0}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\eta_{1}^{i}\theta_{1}^{i}\big{(}f_{i}(x_{0}^{i})-f_{i}(x^{\star})\big{)}:=D.$ (41)
-Observe that (39) is of the form $\alpha_{t}^{i}[f_{i}(x_{t}^{i})-f_{i}(x^{\star})]+\sum_{k=1}^{t-1}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]=\sum_{k=1}^{t}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]$.
+Observe that (40) is nonnegative by the definition of $\eta_{t}^{i},$ implying the iterates $\{x_{t}^{i}\}$ are bounded.
+Now, define the set $\mathcal{S}:=\mathrm{conv}(\{x_{0}^{i},x_{1}^{i},\dots\})$, which is bounded as the convex hull of bounded points.
+Therefore, $\nabla f_{i}$ is Lipschitz continuous on $\mathcal{S}$.
+That is, there exist $\hat{L}_{i}$ such that
+$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leqslant\hat{L}_{i}\|x-y\|,\quad\forall x,y\in\mathcal{S}.$
+Also note that (40) is of the form $\alpha_{t}^{i}[f_{i}(x_{t}^{i})-f_{i}(x^{\star})]+\sum_{k=1}^{t-1}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]=\sum_{k=1}^{t}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]$.
 The sum of the coefficients equals:
@@
 Now, notice that by definition of $\eta_{t}^{i},$ we have
-$\eta_{t}^{i}=\frac{\|x_{t}^{i}-x_{t-1}^{i}\|}{2\|\nabla f(x_{t}^{i})-\nabla f(x_{t-1}^{i})\|}\geqslant\frac{1}{2\tilde{L}_{i}}\geqslant\frac{1}{2L},\quad\forall i\in[m].$
-where $\tilde{L}_{i}$ is defined in (10), which by definition is smaller than or equal to the global smoothness $L$.
+$\eta_{t}^{i}=\frac{\|x_{t}^{i}-x_{t-1}^{i}\|}{2\|\nabla f(x_{t}^{i})-\nabla f(x_{t-1}^{i})\|}\geqslant\frac{1}{2\hat{L}_{i}},\quad\forall i\in[m].$
 Also, notice that $\eta_{1}^{i}\theta_{1}^{i}=\frac{(\eta_{1}^{i})^{2}}{\eta_{0}^{i}}=(\eta_{1}^{i})^{2},$ assuming for simplicity that $\eta_{0}^{i}=1.$ Thus, going back to (44), we have
-$S_{t}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}$ $\geqslant\sum_{k=1}^{t}\frac{1}{2\tilde{L}_{i}}+\frac{1}{4\tilde{L}_{i}^{2}}$
-$=\frac{t}{2L_{i}}+\frac{1}{4L_{i}^{2}}\geqslant\frac{t}{2L}+\frac{1}{4L^{2}}.$
-Finally, we have
+$S_{t}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}$ $\geqslant\sum_{k=1}^{t}\frac{1}{2\hat{L}_{i}}+\frac{1}{4\hat{L}_{i}^{2}}$
+$=\frac{t}{2\hat{L}_{i}}+\frac{1}{4\hat{L}_{i}^{2}}\geqslant\frac{t}{2\hat{L}}+\frac{1}{4\hat{L}^{2}},$
+where $\hat{L}=\max_{i}\hat{L}_{i}.$ Finally, we have
 $\frac{D}{2}$ $\geqslant\frac{1}{m}\sum_{i=1}^{m}S_{t}^{i}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}$
-$\geqslant\frac{1}{m}\sum_{i=1}^{m}\Big{(}\frac{t}{2L}+\frac{1}{4L^{2}}\Big{)}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)},$
+$\geqslant\frac{1}{m}\sum_{i=1}^{m}\Big{(}\frac{t}{2\hat{L}}+\frac{1}{4\hat{L}^{2}}\Big{)}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)},$
 which implies
-$\frac{2DL^{2}}{2Lt+1}\geqslant\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}.$
-A.3 Proof of adaptive gradient descent in nonconvex case: centralized
-For the centralized version, we use the original step size from Malitsky and Mishchenko (2020) as in (4).
-Theorem 6.
-Running gradient descent with step size in (4) for nonconvex $f$, it holds that
-$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(x_{t})\|^{2}\leqslant\frac{8\tilde{L}}{3T}[f(x_{0})-f^{\star}],$
-[…]
-$\|\nabla f(x)-\nabla f(y)\|\leqslant\tilde{L}\|x-y\|\quad\forall x,y\in\mathcal{C}$
-Proof.
-Using Lemma 2, where with slight abuse of notation we interpret $f_{i}$ as $f$ and $\eta_{t}^{i}$ as $\eta_{t},$ we have
-$f(x_{t+1})$ $\leqslant f(x_{t})+\langle\nabla f(x_{t}),x_{t+1}-x_{t}\rangle+\frac{1}{4\eta_{t}}\|x_{t+1}-x_{t}\|^{2}$
-$=f(x_{t})+\langle\nabla f(x_{t}),-\eta_{t}\nabla f(x_{t})\rangle+\frac{1}{4\eta_{t}}\|x_{t+1}-x_{t}\|^{2}$
-$=f(x_{t})-\eta_{t}\|\nabla f(x_{t})\|^{2}+\frac{\eta_{t}}{4}\|\nabla f(x_{t})\|^{2}$
-$=f(x_{t})-\frac{3\eta_{t}}{4}\|\nabla f(x_{t})\|^{2}$
-Rearranging, we have
-$\|\nabla f(x_{t})\|^{2}\leqslant\frac{4}{3\eta_{t}}[f(x_{t})-f^{\star}+f^{\star}-f(x_{t-1})]$
-Unfolding for $T$ iterations and using telescoping, we get
-$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(x_{t})\|^{2}\leqslant\frac{4}{3T\eta_{t}}[f(x_{0})-f^{\star}]\leqslant\frac{8\tilde{L}}{3T}[f(x_{0})-f^{\star}]$
-as desired. ∎
+$\frac{2D\hat{L}^{2}}{2\hat{L}t+1}\geqslant\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}.$
 Appendix B Additional Experiments and Plots
@@
 Figure 4: Effect of using different $\delta$ in the second condition of the step size of $\Delta$-SGD.
-B.2 Illustration of the degree of non-iidness
-As mentioned in the main text, the level of "non-iidness" is controlled using latent Dirichlet allocation (LDA) applied to the labels.
-This approach is based on Hsu et al.
-(2019) and involves assigning each client a multinomial distribution over the labels.
-This distribution determines the labels from which the client’s training examples are drawn.
-Specifically, the local training examples for each client are sampled from a categorical distribution over $N$ classes, parameterized by a vector $\mathbf{q}$.
-The vector $\mathbf{q}$ is drawn from a Dirichlet distribution with a concentration parameter $\alpha$ times a prior distribution vector $\mathbf{p}$ over $N$ classes.
-To vary the level of "non-iidness," different values of $\alpha$ are considered: $0.01,0.1$, and $1$.
-A larger $\alpha$ corresponds to settings that are closer to i.i.d.
-scenarios, while smaller values introduce more diversity and non-iidness among the clients, as visualized in Figure 5.
-As can be seen, for $\alpha=1,$ each client (each row on $y$-axis) have fairly diverse classes, similarly to the i.i.d.
-case; as we decrease $\alpha,$ the number of classes that each client has access to decreases.
-With $\alpha=0.01,$ there are man y clients who only have access to a single or a couple classes.
-Figure 5: Illustration of the degree of heterogeneity induced by using different concentration parameter $\alpha$ for Dirichlet distribution, for CIFAR-10 dataset (10 colors) and 100 clients (100 rows on $y$-axis).
-B.3 Effect of different number of local iterations
-In this subsection, we show the effect of different number of local epochs performed per client, while fixing the experimental setup to be the classification of CIFAR-100 datset, trained with a ResNet-50.
+B.2 Effect of different number of local iterations
+In this subsection, we show the effect of the different number of local epochs, while fixing the experimental setup to be the classification of CIFAR-100 dataset, trained with a ResNet-50.
 We show for all three cases of the Dirichlet concentration parameter $\alpha\in\{0.01,0.1,1\}$.
 The result is shown in Figure 5.
-As mentioned in the main text, instead of performing local iterations, we instead perform local epochs $E,$ similarly to Reddi et al.
-(2021); Li et al.
-(2020).
+As mentioned in the main text, instead of performing local iterations, we instead perform local epochs $E,$ similarly to (Reddi et al., 2021; Li et al., 2020).
 All the results in Table 1 use $E=1,$ and in Figure 5, we show the results for $E\in\{1,2,3\}$.
 Note that in terms of $E$, the considered numbers might be small difference, but in terms of actual local gradient steps, they differ significantly.
-As can be seen, using higher $E$ leads to slightly faster convergence in all cases of $\alpha.$ However, the speed up seems marginal compared to the amount of extra computation time it requires (i.e., using $E=2$ takes roughly twice more total wall-clock time than using $E=1$).
+As can be seen, using higher $E$ leads to slightly faster convergence in all cases of $\alpha.$ However, the speed-up seems marginal compared to the amount of extra computation time it requires (i.e., using $E=2$ takes roughly twice more total wall-clock time than using $E=1$).
 Put differently, $\Delta$-SGD performs well with only using a single local epoch per client ($E=1$), and still can achieve great performance.
-Figure 6:
-B.4 Additional experiments using different number of local data per client
+Figure 5: Effect of the different number of local epochs.
+B.3 Additional experiments using different number of local data per client
+Non-iidness Optimizer Dataset / Model
+$\text{Dir}(\alpha\cdot\mathbf{p})$ CIFAR-10 CIFAR-100
+ResNet-18 ResNet-50
+$\alpha=0.1$ SGD 80.7↓(0.0) 53.5↓(4.0)
+SGD ($\downarrow$)↓ ) 78.8↓(1.9) 53.6↓(3.9)
+SGDM 75.0↓(5.7) 53.9↓(3.6)
+SGDM ($\downarrow$)↓ ) 66.6↓(14.1) 53.1↓(4.4)
+Adam 79.9↓(0.8) 51.1↓(6.4)
+Adagrad 79.3↓(1.4) 44.5↓(13.0)
+SPS 64.4↓(16.3) 37.2↓(20.3)
+$\Delta$-SGD 80.4↓(0.3) 57.5↓(0.0)
+Table 3: Experimental results using different number of local data per client.
+Performance difference within 0.5% of the best result for each task are shown in bold.
+Subscripts↓(x.x) is the performance difference from the best result and is highlighted in pink when the difference is bigger than 2%.
+The symbol ($\downarrow$) appended to SGD and SGDM indicates step-wise learning rate decay, where the step sizes are divided by $10$ after 50%, and another by $10$ after 75% of the total rounds.
 In this subsection, we provide additional experimental results, where the size of the local dataset each client has differs.
@@
 We keep all other settings the same.
-Then, in the server aggregation step (line 13 of Algorithm 1), we compute the weighted average of the local parameters, instead of the plain average, as suggested in McMahan et al.
-(2017), and used for instance in Reddi et al.
-(2021).
-We experiment in two settings: CIFAR-10 dataset classification trained with a ResNet-18 He et al.
-(2016), with Dirichlet concentration parameter $\alpha=0.1,$ which was the setting where the algorithms were fined-tuned using grid search; please see details in Section 4.
+Then, in the server aggregation step (line 13 of Algorithm 1), we compute the weighted average of the local parameters, instead of the plain average, as suggested in (McMahan et al., 2017), and used for instance in (Reddi et al., 2021).
+We experiment in two settings: CIFAR-10 dataset classification trained with a ResNet-18 (He et al., 2016), with Dirichlet concentration parameter $\alpha=0.1,$ which was the setting where the algorithms were fined-tuned using grid search; please see details in Section 4.
 We also provide results for CIFAR-100 dataset classification trained with a ResNet-50, with the same $\alpha.$
@@
 For CIFAR-100 classification trained with ResNet-50, $\Delta$-SGD outperforms all other methods with large margin.
+B.4 Additional experiments using FedAdam
+As mentioned in Section 2 of the main text, since our work focuses on client adaptivity, our method can seamlessly be combined with server-side adaptivity like FedAdam (Reddi et al., 2021).
+Here, we provide additional experiments using the FedAdam server optimizer.
+A big downside of the adaptive server-side approach like FedAdam is that it adds at least one more hyperparameter to tune: the step size for the global optimizer, set aside the two “momentum” parameters ($\beta_{1}$ and $\beta_{2}$) in the case of Adam.
+This necessity to tune additional parameters is quite orthogonal to the approach we take, where we try to remove the need for step size tuning.
+Indeed, as can be seen in Reddi et al.
+(2021)[Appendix D.2], 9 different server-side global optimizer step size is grid-searched, and Reddi et al.
+(2021)[Table 8] shows the best performing server learning rate differs for each task.
+That being said, we simply used the default step size and momentum values of Adam (from Pytorch), to showcase that our method can seamlessly combined with FedAdam.
+Due to time constraints, we tested up to CIFAR-100 trained with a ResNet-18, so that all datasets are covered.
+The result can be found below.
+For three out of four cases, $\Delta$-SGD equipped with Adam server-side optimizer achieves the best accuracy with large margin.
+For the last case, CIFAR-100 trained with a ResNet-18, SGD with decaying step size achieves noticeably better accuracy than $\Delta$-SGD; however, we refer the readers to Table 1 where for the same setting, $\Delta$-SGD equipped with simple averaging achieves $61.1\%$ accuracy, without any tuning.
+Additional experiments using FedAdam (Reddi et al., 2021)
 Non-iidness Optimizer Dataset / Model
-$\text{Dir}(\alpha\cdot\mathbf{p})$ CIFAR-10 CIFAR-100
-ResNet-18 ResNet-50
-$\alpha=1$ SGD 80.7${}_{\downarrow(0.0)}$ 53.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.0)}}$
-SGD ($\downarrow$)↓ ) 78.8${}_{\downarrow(1.9)}$ 53.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.9)}}$
-SGDM ${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.7)}}$ 53.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.6)}}$
-SGDM ($\downarrow$)↓ ) 66.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.1)}}$ 53.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.4)}}$
-Adam 79.9${}_{\downarrow(0.8)}$ 51.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.4)}}$
-Adagrad 79.3${}_{\downarrow(1.4)}$ 44.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.0)}}$
-SPS 64.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.3)}}$ 37.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(20.3)}}$
-$\Delta$-SGD 80.4${}_{\downarrow(0.3)}$ 57.5${}_{\downarrow(0.0)}$
-Table 2: Experimental results using different number of local data per client.
-Performance difference within 0.5% of the best result for each task are shown in bold.
-Subscripts${}_{\downarrow(x.x)}$ is the performance difference from the best result and is highlighted in pink when the difference is bigger than 2%.
+$\text{Dir}(\alpha\cdot\mathbf{p})$ MNIST FMNIST CIFAR-10 CIFAR-100
+CNN CNN ResNet-18 ResNet-18
+$\alpha=0.1$ SGD $97.3_{\downarrow(0.6)}$ $83.7_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ $52.0_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.8)}}$ $46.7_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.5)}}$
+SGD ($\downarrow$)↓ ) $96.4_{\downarrow(1.4)}$ $80.9_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.9)}}$ $49.1_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.7)}}$ ${\color[rgb]{0.13,0.55,0.13}\textbf{49.2}}_{\downarrow(0.0)}$
+SGDM ${\color[rgb]{0.13,0.55,0.13}\textbf{97.5}}_{\downarrow(0.4)}$ $84.6_{\downarrow(1.2)}$ $53.7_{{\color[rgb]{0.96,0.29,0.54}\downarrow(10.1)}}$ $13.3_{{\color[rgb]{0.96,0.29,0.54}\downarrow(35.9)}}$
+SGDM ($\downarrow$)↓ ) $96.4_{\downarrow(1.5)}$ $81.8_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.0)}}$ $53.3_{{\color[rgb]{0.96,0.29,0.54}\downarrow(10.5)}}$ $16.8_{{\color[rgb]{0.96,0.29,0.54}\downarrow(32.4)}}$
+Adam $96.4_{\downarrow(1.5)}$ $81.5_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.3)}}$ $27.8_{{\color[rgb]{0.96,0.29,0.54}\downarrow(36.0)}}$ $38.3_{{\color[rgb]{0.96,0.29,0.54}\downarrow(10.9)}}$
+Adagrad $95.7_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.2)}}$ $82.1_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.7)}}$ $10.4_{{\color[rgb]{0.96,0.29,0.54}\downarrow(53.4)}}$ $1.0_{{\color[rgb]{0.96,0.29,0.54}\downarrow(48.2)}}$
+SPS $96.6_{\downarrow(1.3)}$ $85.0_{\downarrow(0.8)}$ $21.6_{{\color[rgb]{0.96,0.29,0.54}\downarrow(42.2)}}$ $1.6_{{\color[rgb]{0.96,0.29,0.54}\downarrow(47.6)}}$
+$\Delta$-SGD ${\color[rgb]{0.13,0.55,0.13}\textbf{97.9}}_{\downarrow(0.0)}$ ${\color[rgb]{0.13,0.55,0.13}\textbf{85.8}}_{\downarrow(0.0)}$ ${\color[rgb]{0.13,0.55,0.13}\textbf{63.8}}_{\downarrow(0.0)}$ $41.9_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.3)}}$
+Table 4: Additional experiments using FedAdam (Reddi et al., 2021).
+B.5 Additional experiments using MOON
+Here, we provide additional experimental results using MOON (Li et al., 2021), which utilizes model-contrastive learning to handle heterogeneous data.
+In this sense, the goal of MOON is very similar to that of FedProx (Li et al., 2020); yet, due to the model-contrastive learning nature, it can incur bigger memory overhead compared to FedProx, as one needs to keep track of the previous model(s) to compute the representation of each local batch from the local model of last round, as well as the global model.
+As such, we only ran using MOON for CIFAR-10 classification using a ResNet-18; for the CIFAR-100 classification using a ResNet-50, our computing resource ran out of memory to run the same configuration we ran in the main text.
+The results are summarized in Table 5.
+As can be seen, even ignoring the fact that MOON utilizes more memory, MOON does not help for most of the optimizers–it actually often results in worse final accuracies than the corresponding results in Table 1 from the main text.
+Non-iidness Optimizer Dataset / Model
+$\text{Dir}(\alpha\cdot\mathbf{p})$ CIFAR-10
+ResNet-18
+$\alpha=0.1$ SGD 78.2↓(4.9)
+SGD ($\downarrow$)↓ ) 74.2↓(8.9)
+SGDM 76.4↓(6.7)
+SGDM ($\downarrow$)↓ ) 75.5↓(14.1)
+Adam 82.4↓(0.6)
+Adagrad 81.3↓(1.8)
+SPS 9.57↓(73.5)
+$\Delta$-SGD 83.1↓(0.0)
+Table 5: Experimental results using MOON loss function.
+Performance differences within 0.5% of the best result for each task are shown in bold.
+Subscripts↓(x.x) is the performance difference from the best result and is highlighted in pink when the difference is bigger than 2%.
 The symbol ($\downarrow$) appended to SGD and SGDM indicates step-wise learning rate decay, where the step sizes are divided by $10$ after 50%, and another by $10$ after 75% of the total rounds.
+B.6 Additional experiments using FedProx
+Here, we provide the complete result of Table 4(b), conducting the experiment using FedProx loss function for the most heterogeneous case ($\alpha=0.01$) for all datasets and architecture considered.
+The result can be found below, where $\Delta$-SGD achieves the best accuracy in three out of five cases (FMNIST trained with a CNN, CIFAR-10 trained with a ResNet-18, and CIFAR-100 with ResNet-50); in the remaining two cases (MNIST trained with a CNN and CIFAR-100 trained with a ResNet-18), $\Delta$- SGD achieves second-best accuracy with close margin to the best.
+Additional experiments using FedProx loss function (Li et al., 2020)
+Non-iidness Optimizer Dataset / Model
+$\text{Dir}(\alpha\cdot\mathbf{p})$ MNIST FMNIST CIFAR-10 CIFAR-100 CIFAR-100
+CNN CNN ResNet-18 ResNet-18 ResNet-50
+$\alpha=0.01$ SGD 95.7↓(1.4) 79.0↓(1.6) 20.0↓(13.8) 30.3↓(0.0) 25.2↓(5.9)
+SGD ($\downarrow$)↓ ) 97.2↓(0.0) 79.3↓(1.3) 31.3↓(2.5) 29.5↓(0.8) 20.2↓(10.8)
+SGDM 73.8↓(23.4) 72.8↓(7.8) 29.3↓(4.4) 22.9↓(7.5) 23.8↓(7.2)
+SGDM ($\downarrow$)↓ ) 81.2↓(16.0) 78.0↓(2.6) 25.3↓(8.5) 19.8↓(10.5) 15.0↓(16.0)
+Adam 82.3↓(14.8) 65.6↓(15.0) 28.1↓(5.7) 24.8↓(5.6) 22.6↓(8.4)
+Adagrad 76.3↓(20.9) 51.2↓(29.4) 19.3↓(14.5) 12.9↓(17.4) 4.1↓(26.9)
+SPS 85.5↓(11.7) 62.1↓(18.45) 27.6↓(6.2) 21.3↓(9.0) 16.5↓(14.5)
+$\Delta$-SGD 96.9↓(0.26) 80.6↓(0.0) 33.8↓(0.0) 29.7↓(0.6) 31.0↓(0.0)
+Table 6: Additional experiments using FedProx loss function.
+B.7 Additional plot with standard deviation (3 random seeds)
+In this section, we repeat for three independent trials (using three random seeds) for two tasks: CIFAR-10 classification trained with a ResNet-18, and FMNIST classification trained with a shallow CNN, both with Dirichlet $\alpha=0.1,$.
+The aim is to study the robustness of $\Delta$-SGD in comparison with other client optimizers.
+We remind the readers that CIFAR-10 classification trained with a ResNet-18 with Dirichlet $\alpha=0.1$ is where we performed grid-search for each client optimizer, and thus each are using the best step size that we tried.
+Then, for FMNIST classification, the same step sizes from the previous task are used.
+The result can be found below.
+We see that $\Delta$-SGD not only achieves the best average test accuracy, but also achieves very small standard deviation.
+Critically, none of the other client optimizers achieve good performance in both settings, other than the exception of $\Delta$-SGD.
+Figure 6: Additional plots for CIRAR-10 classification trained with a ResNet-18, FMNIST classification trained with a CNN, and CIFAR-100 classification trained with a ResNet-50, repeated for three independent trials.
+The average is plotted, with shaded area indicating the standard deviation.
+Non-iidness Optimizer Dataset / Model
+$\text{Dir}(\alpha\cdot\mathbf{p})$ CIFAR-10/ResNet-18 FMNIST/CNN CIFAR-100/ResNet-50
+Average Std. Dev. Average Std. Dev. Average Std. Dev.
+$\alpha=0.1$ SGD 73.96↓(9.93) 0.0263 83.43↓(1.78) 0.0178 49.44↓(14.58) 0.0376
+SGD ($\downarrow$)↓ ) 67.79↓(16.10) 0.1550 77.83↓(7.38) 0.0815 40.26↓(23.76) 0.1687
+SGDM 77.59↓(6.30) 0.0142 83.27↓(1.94) 0.0186 53.80↓(10.22) 0.0149
+SGDM ($\downarrow$)↓ ) 72.11↓(11.78) 0.0611 81.36↓(3.85) 0.0606 46.77↓(17.25) 0.0762
+Adam 83.24↓(0.65) 0.0126 79.98↓(5.23) 0.0228 58.89↓(5.13) 0.0098
+Adagrad 83.86↓(0.03) 0.0048 53.06↓(32.15) 0.0487 49.93↓(14.09) 0.0177
+SPS 68.38↓(15.51) 0.0192 84.59↓(0.62) 0.0154 43.16↓(20.86) 0.0072
+$\Delta$-SGD 83.89↓(0.00) 0.0052 85.21↓(0.00) 0.0152 64.02↓(0.00) 0.0051
+Table 7: CIRAR-10 classification trained with a ResNet-18 and FMNIST classification trained with a CNN, and CIFAR-100 classification trained with a ResNet-50, repeated for three independent trials.
+The average among three trials and the standard deviation for each client optimizer are reported.
+B.8 Step size plot for $\Delta$-SGD
+Here, we visualize the step size conditions for $\Delta$-SGD to see how the proposed step size looks like in practice. We only plot the first 300 epochs, as otherwise the plot gets quite messy.
+From Figure 7, we can see that both conditions for $\eta_{t,k}^{i}$ are indeed necessary.
+The first condition, plotted in green, approximates the local smoothness of $f_{i},$ but can get quite oscillatory.
+The second condition, plotted in blue, effectively restricts the first condition from taking too large values.
+Figure 7: Step size conditions for $\Delta$-SGD.
+In green, we plot the first condition, and in blue, we plot the second condition for $\eta_{t,k}^{i}$, which is plotted in red color taking the minimum of the two conditions.
+B.9 Illustration of the degree of heterogeneity
+As mentioned in the main text, the level of “non-iidness” is controlled using latent Dirichlet allocation (LDA) applied to the labels.
+This approach is based on (Hsu et al., 2019) and involves assigning each client a multinomial distribution over the labels.
+This distribution determines the labels from which the client’s training examples are drawn.
+Specifically, the local training examples for each client are sampled from a categorical distribution over $N$ classes, parameterized by a vector $\mathbf{q}$.
+The vector $\mathbf{q}$ is drawn from a Dirichlet distribution with a concentration parameter $\alpha$ times a prior distribution vector $\mathbf{p}$ over $N$ classes.
+To vary the level of “non-iidness,” different values of $\alpha$ are considered: $0.01,0.1$, and $1$.
+A larger $\alpha$ corresponds to settings that are closer to i.i.d.
+scenarios, while smaller values introduce more diversity and non-iidness among the clients, as visualized in Figure 8.
+As can be seen, for $\alpha=1,$ each client (each row on $y$-axis) have fairly diverse classes, similarly to the i.i.d.
+case; as we decrease $\alpha,$ the number of classes that each client has access to decreases.
+With $\alpha=0.01,$ there are man y clients who only have access to a single or a couple classes.
+Figure 8: Illustration of the degree of heterogeneity induced by using different concentration parameter $\alpha$ for Dirichlet distribution, for CIFAR-10 dataset (10 colors) and 100 clients (100 rows on $y$-axis).